### v1

# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling

Abstract
Transformers have become the cornerstone of modern large-scale language models; however, their dependence on softmax attention poses a major computational bottleneck, particularly in long-context settings. In this work, rather than following prevalent approaches such as linear attention (or SSMs) and local attention, we introduce an intermediate design called RAT between recurrence and attention mechanisms. It partitions the input into chunks, applies a simple linear recurrence within each chunk to capture local dependencies, and then performs softmax attention across chunks to model long-range interactions. By adjusting the size of the chunk, RAT enables flexible trade-offs, combining the strengths of RNN and attention. Empirically, with a chunk size of 16, the RAT layer achieves a 7×7$\times$7 × improvement in training speed with 100K token sequences and 9×9$\times$9 × in generation at 4K sequence length, while maintaining similar or sometimes even better accuracy compared to standard attention. We demonstrate this by training 1.3B parameter models from scratch and performing large-scale evaluations, including short- and long-context benchmarks, as well as supervised fine-tuning (SFT). We further propose a hybrid architecture that interleaves RAT with local attention. By combining efficient long-range modeling with strong local interactions, this hybrid design not only improves inference speed and reduces cache memory usage compared to attention, but also consistently enhances performance, for example, achieving an average 1 point gain in commonsense reasoning tasks, up to 4 points on code tasks, and a 1 point Rouge-L increase in a summarization SFT task. Code is available at https://github.com/CLAIRE-Labo/RAT.
1 Introduction
Language modeling has long been dominated by Transformer-based architectures because of their remarkable performance across a wide range of tasks. However, their reliance on full self-attention leads to quadratic complexity with respect to sequence length, which limits scalability in long-context processing. This has motivated many efforts to design more efficient attention mechanisms. Some approaches focus on local attention [1, 2, 3], and use sparse attention or recurrence to capture long-range information. Another line of work investigates the duality between linear RNNs and attention [4, 5, 6, 7], by enforcing strict linearity constraints and introducing the notion of state expansion, which represents recurrent states as matrices rather than vectors, and shifting from per-dimension to per-head gating.
In contrast to them, we propose an alternative architecture. We argue that overusing attention in short-range contexts underutilizes its strengths and that such local processing can instead be handled more efficiently by lightweight structures such as recurrences. Based on this insight, we explore a design that lies between RNNs and attention. This design also avoids strict linearity assumptions on either side and adopts the classical recurrence form with per-dimension gating and no state expansion.
Figure 1: Comparison of three architectures: The circles represent the token representations and the arrows indicate the information flow for the 5th token. (a) Attention: Each token retains its representation and has full access to all previous tokens. (b) RNN: The information is progressively compressed along the sequence, and the current token only accesses the last hidden state. (c) RAT(L=2): Intra-chunk recurrence compresses local information, while inter-chunk attention allows direct access to previous chunks. Note that (a) and (b) can be viewed as special cases of RAT with $L=1$ and $L=T$, respectively.
Figure 2: RAT structure. The symbol $\sigma(\cdot)$ further applies a sigmoid function to the gates. Note that we do not use extra convolution or normalization in our design.
We propose a RAT layer, which is a simple but effective temporal mixing mechanism. We divide long sequences into chunks: within each chunk, we apply recurrence for local modeling, and use softmax-based attention across chunks to access distant information directly (see Fig. 1). Recurrence efficiently captures short-range dependencies without suffering from the memory degradation issues common in long sequences, while attention over chunk-level representations enables long-range retrieval. Adjusting the size of the fragment $L$, RAT interpolates between attention (when $L=1$) and RNN (when $L=T$). To make training more efficient, we adopt a lightweight gated linear recurrence [8]. We further explore a hybrid design that interleaves RAT with sliding-window attention layers [1], allowing the two mechanisms to complement each other while remaining efficient.
In terms of efficiency, our implementation relies solely on PyTorch high-level operators without any custom CUDA kernels, yet already achieves a substantial speed-up. With a chunk size of $L=16$, the single layer latency to generate the token at position 4096 is up to 9×9$\times$9 × faster than that of vanilla attention. And as shown in Table 1, the maximum throughput of the full model improves by up to 10×10$\times$10 × over attention, and interleaving with local attention can also yield a 4×4$\times$4 × gain.
We pretrained models with different chunk sizes at the 1.3B scale using a 4K context length, and found their perplexity (PPL) falls between that of RNNs and attention. We then evaluated them in three settings: (1) zero-shot commonsense reasoning with short contexts, (2) long-context understanding on LongBench [9], and (3) SFT on two question answering (QA) and two summarization tasks. RAT with $L=16$ performs comparably to attention on most benchmarks and even surpasses it on some LongBench tasks. When interleaved with local attention, it achieves the strongest overall results across all variants while maintaining high efficiency. Our main contributions are as follows:
  1. 1.
    We propose the RAT layer as an intermediate design between RNN and attention for temporal mixing. We explore its detailed design and efficient implementation across three key scenarios, and observe significant speed.
  2. 2.
    We pretrain models at the 1.3B scale and evaluate them on 6 short-context, 14 long-context, and 4 SFT tasks, showing that RAT(L=16) performs competitively with attention.
  3. 3.
    We further investigate a hybrid variant that interleaves RAT with sliding-window attention, which yields the best overall performance—for example, +1 accuracy on commonsense reasoning tasks, +4 on code completion, +4 on a challenging QA SFT task, and +1 on a challenging summarization task—compared to full attention or its local variant.
Table 1: Representative results for 1.3B models across pretraining, direct evaluation, and SFT. -SWA denotes interleaving with sliding-window attention (SWA) (window size 1024). Maximum throughput is measured by generating 1024 tokens given a prompt of 3072 tokens on a H100 GPU in GH200 system. See Sec. 3 for details.
Model Throughput Pretrain Direct Evaluation SFT Val. CSR SQA Summ Code NQA1 QMSum token/sec PPL Avg. acc Avg. F1 Avg. Rouge-L Avg. EditSum F1 Rouge-L Attention 3052 7.61 56.7 18.2 19.5 23.9 61.3 23.4 RAT(L=16) 31170 7.67 55.6 19.6 20.2 17.4 60.8 23.3 Attention-SWA 4605 7.61 56.4 17.4 19.4 21.7 63.3 23.4 RAT(L=16)-SWA 13582 7.57 57.8 18.8 19.5 28.2 63.2 24.6
2 Method
2.1 Attention in temporal mixing
In softmax attention mechanisms, each token has access to all preceding tokens through a learned token-specific weighted aggregation:
${\bm{y}}_{t}=f({\bm{q}}_{t}{\bm{K}}^{\top}_{:}){\bm{V}}_{:}$ (1)
$f(\cdot)$ denotes causal masking and the softmax function, and $f({\bm{q}}_{t}{\bm{K}}^{\top}_{:})$ are the attention weights used to aggregate ${\bm{V}}_{:}$. We use Python-style indexing in ${\bm{K}}^{\top}_{:},{\bm{V}}_{:}$ to highlight that each query is concerned with all keys and values. Although attention is widely considered a crucial component in state-of-the-art architectures for sequence modeling, its full-token access incurs a computational complexity of $\mathcal{O}(T^{2})$, where $T$ is the length of the sequence. This quadratic growth in FLOPs can lead to substantial inefficiency during training and inference. We view attention as a slow but lossless module in the mixing of information between tokens and aims to combine it with a fast but lossy component.
2.2 Recurrence in temporal mixing
RNNs exploit sequential dependencies in the data [10, 11]. As the model processes the sequence, it maintains a summary of past information in its hidden state, allowing each step to be computed solely from the previous state and the current input. This enables efficient inference and low FLOPs, in contrast to the full-sequence access required by attention. To further improve training efficiency, recent work has explored linear RNN variants [12, 13]. For simplicity, we adopt the design proposed in Feng et al. [8], which is similar to that in Smith et al. [14]. Using diagonal matrices and removing nonlinearities, the recurrence reduces to an EMA-like gating behavior, which can be efficiently implemented with the parallel scan algorithm [14]:
$\tilde{{\bm{v}}}_{t}$ $={\bm{g}}_{t}\odot\tilde{{\bm{v}}}_{t-1}+(1-{\bm{g}}_{t})\odot{\bm{v}}_{t},$ (2)
${\bm{y}}_{t}$ $={\bm{z}}_{t}\odot\tilde{{\bm{v}}}_{t},$
where ${\bm{g}}_{t}$ and ${\bm{z}}_{t}\in\mathbb{R}^{D}$ denote the forget and output gates, respectively, computed via linear projections of the input followed by a sigmoid activation. Here, $D$ refers to the dimension of the model.
We find this simple design sufficient to achieve strong performance for short sequences. However, compressing the entire history to the current state, even with more complex recurrences, can lead to memory degradation over time [15]. As a result, RNN may struggle to retain information from distant time steps, which poses challenges for long-context tasks. Thus, in this paper, we place RNN as the fast yet lossy token mixing module.
2.3 RAT: a design between RNN and attention
Motivated by the complementary strengths of RNN and attention, we propose an intermediate design that combines their advantages: the speed of RNN and the global access of attention. Since RNN are effective for short sequences, while attention excels at modeling long-range dependencies, we reinterpret the input as a sequence of shorter chunks of sequences. A recurrent module is applied within each chunk to capture local dependencies, followed by cross-chunk attention to model global interactions efficiently, as shown in Fig. 1.
We divide a sequence of length $T$ into $C$ chunks of length $L$, such that $T=C\cdot L$. A token originally at position $t$ is re-indexed as $(c,l)$, where $c$ denotes the chunk index and $l$ the position within the chunk. Within each chunk, a forget gate ${\bm{g}}_{c,l}$ is used to recurrently aggregate the value ${\bm{v}}_{c,l}$ and key ${\bm{k}}_{c,l}$ vectors, yielding updated representations $\tilde{{\bm{v}}}_{c,l}$ and $\tilde{{\bm{k}}}_{c,l}$. For each query ${\bm{q}}_{c,l}$, we compute attention on the final key vectors $\tilde{{\bm{K}}}_{:,-1}$ from all preceding chunks, as well as the current chunk’s key $\tilde{{\bm{k}}}_{c,l}$. The causal masking function $f(\cdot)$ restricts attention to the chunks before it, followed by a softmax operation. Finally, an output gate is applied to produce the output:
$\tilde{{\bm{v}}}_{c,l}={\bm{g}}_{c,l}\odot\tilde{{\bm{v}}}_{c,l-1}+(1-{\bm{g}}_{c,l})\odot{\bm{v}}_{c,l}$    (Intra-chunk RNN) (3)
$\tilde{{\bm{k}}}_{c,l}={\bm{g}}_{c,l}\odot\tilde{{\bm{k}}}_{c,l-1}+(1-{\bm{g}}_{c,l})\odot{\bm{k}}_{c,l}$    (Intra-chunk RNN)
${\bm{y}}_{c,l}=f([{\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top};{\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top}])[\tilde{{\bm{V}}}_{:,-1};\tilde{{\bm{v}}}_{c,l}]$    (Inter-chunk attention)
${\bm{y}}_{c,l}={\bm{z}}_{c,l}\odot{\bm{y}}_{c,l}$
The overall computation flow is illustrated in Fig. 2. Benefiting from the short length of each chunk, the RNN is both effective and computationally efficient, since its hidden state has sufficient capacity to encode short-range context without discarding much information. The attention applied across chunks then enables access to distant positions without requiring compression. As a result, the overall design preserves accuracy comparable to that of full softmax attention, while reducing computational cost, as inter-chunk attention operates over shorter contexts, with attention length inversely proportional to the chunk size.
We position this design as an intermediate architecture design between RNN and attention, aiming to combine their respective strengths, and refer to it as RAT. By adjusting the size of the chunk $L$, RAT interpolates between the RNN and the attention behaviors: when $L=T$, the model reduces to an RNN, since the entire sequence is treated as a single chunk; when $L=1$, it resembles the attention, since each token resides in its chunk. We also explored a reversed variant that applies attention within chunks and recurrence across chunks. Although this design also allows interpolation, it underperforms RAT with equal FLOPs, likely due to suboptimal allocation of computation to local attention. Further discussion is provided in Appendix C. In the following two paragraphs, we discuss the detailed design choices for RAT.
Parameter allocation
We aim to match the parameter budget of RNN and attention. In addition to the projection of the output, which introduces the parameters $D^{2}$ in all three mechanisms, both attention and RNN use additional parameters $3D^{2}$. These correspond to projections for the query, key, and value in attention, and for the value, the forget gate, and the output gate in RNN. To maintain the exact total parameter count for RAT, which combines both mechanisms, we adopt lightweight designs. Specifically, we consider using low-rank projections for the gates ${\bm{g}}$ and ${\bm{z}}$, or sharing query and key projections across heads. Empirical results in Subsec. 3.3 show that sharing query ${\bm{q}}$ and key ${\bm{k}}$ typically improves performance by enabling more efficient parameter allocation. Importantly, this design does not reduce to single-head attention because the forget gate operates at a per-dimension level and introduces different elements into the gated key $\tilde{{\bm{k}}}$ used for the inter-chunk attention.
Positional encoding
We examine how to encode positional information in RAT. Although RNN captures position through its sequential structure, the attention is usually based on explicit encodings such as RoPE [16]. For RAT, we find that applying RoPE based on chunk indices rather than original token positions yields slightly better performance, and we use cross-chunk RoPE by default, as it may better align with inter-chunk attention. See Subsec. C.2 for a more detailed discussion, including exploration of the NoPE [17] encoding.
2.4 Efficiency
We discuss the efficient implementation of RAT for the three main scenarios in language modeling, and provide pseudocode for training or prefilling in 1 and for generation in 2.
Training or prefilling
In this setting, a sequence of tokens is processed in parallel, requiring the model to support efficient parallel computation. For the intra-chunk recurrent component in Eq. (missing) 3, we use PyTorch’s higher order operator associative scan to implement forward and backward passes with $\mathcal{O}(T)$ FLOPs. Compared to RNN, RAT applies the recurrent operation to both key and value representations, but chunking can reduce the scan depth from $\mathcal{O}(\log{T})$ to $\mathcal{O}(\log{L})$ and increase parallelism. In principle, if the chunk size $L$ is sufficiently small, even non-linear RNNs could be viable in terms of efficiency, which we leave for future work. For inter-chunk attention, each token also attends to its own chunk’s key and value, which vary due to causal masking. To handle this efficiently, we adopt an online softmax formulation [18]: we separately compute $f({\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top})\tilde{{\bm{V}}}_{:,-1}$ and $f({\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top})\tilde{{\bm{v}}}_{c,l}$, then combine them by adjusting the softmax denominator. The first term is computed efficiently and in parallel using PyTorch’s flex attention, which supports custom causal masks and returns the softmax normalization term. The second term is implemented via a simple einsum.
Generation
In this setting, tokens are generated sequentially. The intra-chunk recurrent component can be implemented directly using the recurrence equation, as it only requires a single-step update and does not rely on parallel operators. For inter-chunk attention, standard implementations like flash attention [19] can be used without any modification, since no complex causal masking is required in the generation case.
In summary, the FLOPs per token of RAT is $\mathcal{O}(C\cdot D)$, compared to $\mathcal{O}(D)$ for RNN and $\mathcal{O}(T\cdot D)$ for full attention, where $C$ is the number of chunks, $D$ the model dimension and $T$ the sequence length. Notably, our current implementation achieves significant speed-ups in both training and generation without relying on custom CUDA or Triton kernels, though such kernels could further improve performance. Parallelism is particularly important for models with more than 7B parameters. We believe that RAT is compatible with both tensor and context parallelism. In particular, using a chunk size larger than 16 may reduce the communication overhead of ring-style attention in context parallelism.
3 Experiments
In this section, we present the efficiency (training, prefilling, and generation) and precision (pretraining, evaluation, and SFT) of different models. We focus primarily on the 1.3B model scale with a 4K context window, while additional discussions, including length generalization, 16K context window training, and more evaluation results, are presented in Appendix C.
3.1 Setup
For brevity, we summarize the setup of the 1.3B model with the 4K context window, which is used in most of our experiments, and provide full details in Appendix A.
Model.
We adopt a Transformer architecture that interleaves a token mixing block with a hidden-state mixing block (FFN), each wrapped with residual connections and LayerNorm. We compare variants using different token mixing modules, including RNN, attention, and RAT. Inspired by recent designs that alternate global and sliding-window attention (SWA), we also explore hybrid models. Specifically, Attention-SWA alternates standard attention with SWA (window size 1024) every two layers, while RAT-SWA replaces global attention with RAT.
Efficiency
We benchmark the latency of a single token mixing block, including input and output projections, on the single H100 in the GH200 system with 120GB of memory. For training and prefilling, we report the latency of processing the entire sequence. For generation, which is often the most time-consuming case, we report the latency of generating a batch of tokens at a given position. To ensure a fair comparison with strong baselines, we use flash attention for the attention module and the associative scan operator for RNN. For RAT, we follow our method design as described above. All models are wrapped with torch.compile and evaluated in bfloat16 using torch.cuda.amp.
Accuracy
We pretrain the 1.3B models on a 100B token subset of the FineWeb-Edu dataset [20]. We use a learning rate of 8.0e-4, decayed to 1.0e-6 using cosine annealing, and a global batch size of 2M tokens. Hyperparameters are determined by DeepSeek’s attention-optimized guidelines [21] and are shared across all model variants for fair comparison. Then, we further evaluate our models using two types of benchmarks. The first is a suite of classical commonsense reasoning tasks [22] which are commonly used to assess the reasoning ability of pre-trained models in short contexts. The second is a set of English tasks from LongBench [9], where the average input length for nearly every task exceeds 4,000 tokens, and some tasks have average lengths over 10,000 tokens. LongBench tasks often include specific prompts and instructions, which pretrained-only models struggle to follow reliably. To better assess the model capability, we also compare different architectures after SFT. Specifically, we choose NarrativeQA [23] with two modes, QMSum [24], and WikiSum [25], which vary in sequence length and remain close to the pretraining domain.
3.2 Efficiency study
Fig. 3 presents the efficiency comparison of RNN, RAT, and attention across three scenarios: training, prefilling, and generation. For training and prefilling, when the sequence length is short (e.g., 4096), RAT is slightly slower than attention due to under-utilized GPU parallelism in the flex attention component (with only 256 chunks) and the overhead of associative scan. We believe this can be improved with better CUDA kernel optimization. As the sequence length increases, RAT becomes more efficient: for $L=16$, we observe about 2×2$\times$2 × speedup at 16K, 3×3$\times$3 × at 32K, 4×4$\times$4 × at 64K, and 7×7$\times$7 × at 100K tokens. Additional results in Subsec. B.1 show up to 910×9-10$\times$9 - 10 × at 200K. While RNN slows down beyond 32K, RAT benefits from increased intra-chunk parallelism and retains an advantage over attention. For generations, RAT(L=16) achieves 9×9$\times$9 × speedup at position 4K and around 10×10$\times$10 × for longer sequences. It also reduces the usage of the KV cache, making it much less prone to OOM. RAT(L=64) can be up to 4×4$\times$4 × faster than $L=16$ with larger batch sizes or longer inputs.
Figure 3: Latency results. Latency of a single RNN, RAT, or attention layer (including input and output projections) with model dimension of 2048 under (a) training, (b) prefilling, and (c) generation. (a,b) report full-sequence latency on 200K tokens; (c) reports latency for generating 512 tokens at a fixed position. RAT also reduces KV cache usage, lowering OOM risk. See Appendix B for more results.
3.3 Ablation study
We conduct the ablation study on a 200M-parameter model trained on the book dataset. Allocating more parameters to the intra-chunk RNN gates yields significantly better performance than assigning them to the inter-chunk attention query and key projections, improving the perplexity by 0.4–0.5. Further replacing the original RoPE with a cross-chunk variant, where positions are indexed by chunk index, brings additional improvement, especially at long sequence lengths, with a 0.3 drop in PPL observed.
Figure 4: (a) Ablation study on RAT(L=64). (b) and (c) show pretraining results on 200M and 1.3B models, respectively. RAT lies between RNN and attention in terms of pretraining perplexity.
3.4 Pretraining
We begin our study by examining pretraining perplexity as a foundational measure of architectural performance, showing that RAT lies between attention and RNN. The main results are presented in Fig. 4(b) and (c). In (b), on a 200M model, increasing the size of the chunk leads to higher perplexity. Viewing attention as $L=1$ and RNN as $L=T$, we observe perplexities of 13.26 for attention, 13.42 for RAT(L=64), 13.72 for RAT(L=256), and 14.31 for RNN. Interestingly, increasing the training context from $T=8192$ to $T=32768$ causes a sharp increase in perplexity for RNN, while attention and RAT remain stable. Note that on a 200M model, the PPL may not decrease steadily when going to longer sequences. We then scale up to a 1.3B model trained on web data, which primarily consists of short sequences (often under 2K tokens) concatenated with separator tokens to construct 4K-length contexts. In this regime, we find that the chunk length should be reduced to 16 to match the performance of attention. Once again, increasing the size of the chunks and thus reducing the FLOPs leads to higher perplexity, strengthening the trade-off and intermediate nature of RAT.
3.5 Downstream evaluation
Table 2: We evaluate our model on six commonsense reasoning tasks using the lm-evaluation-harness [22]. These tasks have very short contexts, with each sample containing no more than 300 tokens. Normalized accuracy is reported when available, except for WinoGrande and LAMBADA, which use standard accuracy.
Model ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg.
RNN 38.57 66.46 55.95 39.74 72.69 53.99 54.57
RAT(L=16) 39.33 67.80 56.44 44.40 72.69 53.20 55.64
Attention 37.71 66.67 57.44 47.84 72.91 57.70 56.71
Attention-SWA 37.20 66.20 57.09 47.97 72.74 56.91 56.35
RAT(L=16)-SWA 40.78 68.22 57.85 49.33 73.94 56.91 57.84
Commonsense reasoning
The performance gap between RNN and attention is small on short and relatively simple commonsense reasoning tasks, typically within two points on average. This suggests that RNN can be surprisingly competitive in short-context settings. RAT(L=16) consistently outperforms RNN in most tasks, with particularly strong gains on LAMBADA, highlighting its advantage in modeling token dependencies. However, it performs poorly on WinoGrande compared to attention. One possible explanation is that WinoGrande relies heavily on subtle sentence-level cues and fine-grained referential resolution; moreover, the variance may also stem from the instability of the task itself. Interestingly, interleaving RAT with SWA yields a hybrid model (RAT(L=16)-SWA) that achieves the best overall performance in almost all tasks. This suggests a complementary architecture: using SWA for strong local attention and RAT for long-range dependencies can bring us a model that is not only more efficient but also more accurate than its dense counterpart. In addition, the hybrid version mitigates the weaknesses of RAT in WinoGrande, improving overall robustness. We do not include results from Gated DeltaNet [26], a recent state space model that also interleaves with sliding-window attention, [^1] While our best model outperforms their best model (e.g., 49.33 vs. 48.76 on LAMBADA, 73.94 vs. 72.57 on PIQA, and 57.85 vs. 56.88 on HellaSwag), we do not include their results in our tables. Their approach follows a fundamentally different modeling paradigm—state space modeling—whereas our work focuses on an intermediate design between recurrence and attention. These differing motivations, together with discrepancies in window size (their 2048 vs. our 1024) and training protocols, make direct comparisons less meaningful. as their modeling paradigm and experimental setup differ substantially from ours.
Table 3: Evaluation results on LongBench. For each task, we highlight the top two values to illustrate general trends on instruction-based tasks. To save space, task numbering follows the original LongBench paper; the number-to-task mapping and token counts (measured with the LLaMA2 tokenizer) are provided in Subsec. C.4. Metrics: F1 for QA tasks, Rouge-L for summarization, and EditSum for code tasks. Specifically, accuracy is used for 4-1, F1 for 4-2, and Rouge-L for 4-3.
Model Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
1-1 1-2 1-3 Avg. 2-1 2-2 2-3 Avg. 3-1 3-2 3-3 Avg. 4-1 4-2 4-3 Avg. 6-1 6-2 Avg.
RNN 11.7 8.8 19.5 13.3 9.1 15.1 5.4 9.9 16.9 16.8 16.4 16.7 10.0 17.5 13.6 13.7 14.0 17.6 15.8
RAT(L=64) 13.9 13.0 27.0 18.0 14.9 16.3 4.3 11.8 18.9 17.7 19.9 18.8 17.5 19.2 15.2 17.3 16.2 19.7 18.0
RAT(L=16) 14.5 16.1 28.3 19.6 16.7 18.9 6.3 14.0 24.9 17.5 18.3 20.2 38.5 18.2 18.7 25.1 14.2 20.6 17.4
Attention 12.3 14.0 28.2 18.2 13.4 17.7 4.9 12.0 18.6 18.9 21.0 19.5 39.5 19.6 23.9 27.7 21.3 26.5 23.9
Attention-SWA 13.1 14.6 24.5 17.4 13.7 19.0 6.1 12.9 19.4 17.3 21.4 19.4 47.0 19.6 25.1 30.6 18.7 24.6 21.7
RAT(L=16)-SWA 12.7 15.6 28.2 18.8 14.2 18.1 7.4 13.2 20.1 18.6 19.8 19.5 33.5 18.4 25.6 25.8 26.3 30.1 28.2
LongBench
We present the results of the evaluation for 14 English tasks in LongBench in LABEL:tab:longbench_4k. It is expected that no single model consistently dominates across all tasks, since LongBench employs heavy instruction-style prompts and the tasks span diverse domains. Evaluating pretrained-only models without any fine-tuning on these instruction-based tasks remains challenging; even in the original paper, 7B pretrained-only models perform much worse than tuned ones. Therefore, we focus on general performance trends across tasks, particularly average scores by task category and the top-2 models for each task.
First, RAT(L=16) and its hybrid variant RAT(L=16)-SWA consistently rank among the top two models in nearly all task groups. This is notable given that RAT(L=16) lags behind attention in pretraining perplexity and performs slightly worse in short-context commonsense reasoning. We hypothesize that its advantage on LongBench stems from structural adaptation to very long sequences—for instance, achieving 24.9 on GovReport (3-1) (vs. 18.6 for the attention baseline) and 16.7 on HotpotQA (2-1) (vs. 13.4). Second, unlike the small gap ( 2 points) between RNN and attention on short-context tasks, the difference becomes much larger on LongBench, reaching up to 8 points on code completion. Finally, we note some outliers: RAT-based models underperform on TREC (4-1), a few-shot QA task evaluated via classification accuracy, for reasons not yet fully understood. In addition, RAT(L=16) performs poorly on code completion. However, its hybrid variant RAT(L=16)-SWA significantly improves in this setting, outperforming the second-best model by more than four points, consistent with earlier observations on commonsense reasoning tasks.
3.6 Supervised Fine-tuning
Since LongBench tasks are instruction-based and pretrained-only models often struggle to follow specific prompts, we further assess the performance of SFT on the QA and summarization tasks in the long-context setting, as shown in Table 4.
For NarrativeQA, where passages are extremely long and may not contain the answer within the truncated 4K-token context, we evaluate two settings: (1) summary-only, which uses only the passage summary; and (2) summary+passage, which concatenates the summary and full passage. While the latter provides more context, it also introduces much noise. In both cases, RAT(L=16)-SWA delivers strong performance while remaining efficient. Although the summary and passage contain useful information, we observe that models, particularly the attention module, perform worse in the second setting, with F1 dropping to 33. This highlights the difficulty in identifying and focusing on relevant content in noisy, extended contexts. In this setup, attention can be more easily distracted due to the absence of structural bias, which can lead to flattened attention score distributions [27]. In contrast, models based on RAT appear to be more robust, likely due to their chunk-wise structure. For summarization tasks, RNN performs better than in QA, and RAT(L=16)-SWA achieves the best result on QMSum, surpassing the next-best model by approximately 1 Rouge-L point.
[…]
4 Related work
RNN
Recurrent neural networks (RNNs) have a long history of modeling sequential data [28, 29, 30], but suffer from training instabilities and lack parallelizability, especially when handling long sequences. Various approaches have been proposed to improve their trainability, such as gating mechanisms [10, 11], unitary designs [31, 32], and partially parallelizable variants [33, 34]. Moreover, RNNs face a memory bottleneck as they compress entire sequences into fixed-size hidden-state vectors. Some earlier work addressed this issue by allowing the current step to access earlier time steps directly [35, 36, 37]. For example, Koutnik et al. [36] introduces multiple timescales by updating hidden states only at predefined clock rates, while Ke et al. [37] proposes sparse, learnable connections to past tokens, with weights computed by a small MLP. More recently, state space models [38, 12, 39, 13, 40] have emerged, expanding hidden states from vectors to matrices through state expansion and reduction techniques. They also replace standard matrix multiplications with structured operations (e.g. diagonal matrices) to further enable parallel computation [14]. In contrast to these approaches, and inspired by the success of attention mechanisms, we propose integrating attention directly over the RNN history to allow explicit retrieval of past information. This design remains simple and naturally interpolates between RNN and attention behaviors. By applying inter-chunk attention to retrieve information, we avoid state expansion, which may introduce training instability and typically requires careful normalization or initialization [12, 5].
Attention
While attention mechanisms have proven highly effective, they suffer from slow computation due to their full access to all previous tokens, especially in long sequences. Recently, there has been a growing interest in linear attention [4, 6, 41, 13, 5, 42, 26], which explores the dual formulation of linear RNN and linear attention. These approaches rely on strict linearity constraints and introduce state expansion and per-head gating techniques. In contrast to this duality perspective, we adopt an intermediate view between per-dimension-gated RNNs and softmax-based attention, without imposing hard linearity constraints on either side. The use of linear recurrence in our model is primarily motivated by training efficiency, but it is not strictly required at the design level. Other works have explored attention within local contexts, often motivated by the diagonal structures observed in attention maps, and have introduced recurrence or sparse attention mechanisms to provide access to earlier context [1, 2, 43, 44, 3, 45]. For example, Ma et al. [43] applies attention in chunks after gating the query and key vectors, and Behrouz et al. [3], inspired by cognitive science, focuses attention on the short-range context. In our work, we take a different approach: we use attention to model long-range dependencies, leveraging their strength in directly accessing distant tokens, while applying RNNs to short-range contexts, where they are typically less prone to memory degradation. This can also be seen as a form of attention with resolution, loosely mirroring how humans process information: short-term inputs are integrated into coherent representations, while long-term events are stored as memory anchors that can be selectively retrieved when relevant. We also find that RAT and the attention of sliding windows [1] are structurally complementary. The interleaving of the two leads to further performance improvements while being efficient.
5 Conclusion
This paper proposes the RAT structure, an intermediate architecture between RNN and attention. It segments sequences into chunks, applies RNN within each chunk to capture local dependencies, and employs attention across chunks to model long-range dependencies. We detail its architectural design and investigate its combination with the sliding-window attention mechanism. Experimental results across various settings demonstrate the efficiency and accuracy benefits of RAT, highlighting its potential for future language model development.
Limitations. Due to resource constraints and the exploratory nature of this work, our experiments are limited to models with up to 1.3B parameters. We have not yet scaled RAT to larger language models such as 7B or 14B to confirm our results hold at that scale. Additionally, we did not adopt supervised fine-tuning techniques commonly used in current industry practices, which typically involve substantial data resources and many engineering details. Instead, we followed a more classical approach based on train-test splits to evaluate performance, as our focus lies primarily on architectural design during pretraining. Lastly, RAT may still face length generalization issues from positional encoding, similar to attention. But as shown in Subsec. C.2, this can be significantly reduce by shorter inter-chunk attention spans, and we also explore the use of NoPE.
Appendix A Implementation details
A.1 Algorithm
We provide the pseudocode for the training and prefilling modes of RAT in 1, and the pseudocode for the generation mode in 2.
1def merge_last_token_stable(inter_out, intra_out, inter_lse, intra_lse):
2 # Merge previous chunk’s output and current chunk’s output safely
3 # by adjusting the base of the exponential in softmax
4 max_lse = maximum(inter_lse, intra_lse)
5 inter_lse_exp = exp(inter_lse - max_lse)
6 intra_lse_exp = exp(intra_lse - max_lse)
7 denom = intra_lse_exp + inter_lse_exp
8 intra_adjust = (intra_lse_exp / denom).unsqueeze(-1)
9 inter_adjust = (inter_lse_exp / denom).unsqueeze(-1)
10 return inter_out * inter_adjust + intra_out * intra_adjust
11
12
13def train_or_prefill(z, g, q, k, v, num_chunk, chunk_size, num_head, softmax_scale):
14 """
15 Args:
16 z: (B, T, D), output gate after sigmoid function
17 g: (B, T, D), forget gate after sigmoid function
18 v: (B, T, D), value vectors in attention
19 q, k: (B, T, D), shared query and key vectors in attention
20 Returns:
21 out: (B, T, D)
22 """
23 bs, seq_len, d_model = shape(z)
24 k, v = [rearrange(m, "b (c l) d -> b c l d", c=num_chunk) for m in (k, v)]
25
26 # Intra-chunk RNN: apply associative scan along the "l" dimension
27 intra_v = ascan(g, (1.0 - g) * v)
28 intra_k = ascan(g, (1.0 - g) * k)
29
30 # Rearrange for multi-head attention
31 intra_v = rearrange(intra_v, "b c l (a p) -> b a (c l) p", a=num_head)
32 intra_k = rearrange(intra_k, "b c l (a p) -> b a (c l) p", a=num_head)
33 q = rearrange(q, "b t (a p) -> b a t p", a=num_head)
34
35 # Apply inter-chunk RoPE
36 q, intra_k = apply_rope(q, intra_k)
37
38 # Extract chunk’s end representation for inter-chunk attention
39 chunk_intra_k = intra_k[..., chunk_size - 1::chunk_size, :]
40 chunk_intra_v = intra_v[..., chunk_size - 1::chunk_size, :]
41
42 # Inter-chunk attention using flex_attention
43 # with chunk-level causal mask and returning log-sum-exp scores
44 block_mask = create_block_mask(
45 lambda b, h, q_idx, kv_idx: q_idx // chunk_size > kv_idx,
46 1, 1, q.shape[2], num_chunk
47 )
48 inter_out, inter_lse = flex_attention(
49 q, chunk_intra_k, chunk_intra_v,
50 scale=softmax_scale,
51 block_mask=block_mask,
52 return_lse=True
53 )
54
55 # Compute logits with the current chunk’s k, required separately due to causal masking
56 intra_lse = einsum("batp, batp -> bat", q, intra_k) * softmax_scale
57
58 # Merge outputs from previous chunks and current chunk (due to causal masking)
59 out = merge_last_token_stable(inter_out, intra_v, inter_lse, intra_lse)
60 out = rearrange(out, "b a t p -> b t (a p)", a=num_head)
61 return z * out
Listing 1: Pseudo code for the training or prefilling modes of RAT. We use Pytorch’s flex attention and associative scan for implementation.
1def gen(z, g, q, k, v, chunk_start, num_head, cache):
2 """
3 Args:
4 z: (B, 1, D), output gate after sigmoid
5 g: (B, 1, D), forget gate after sigmoid
6 v: (B, 1, D), value vector
7 q, k: (B, 1, D), shared query and key vectors across heads
8 chunk_start: int, index of current chunk
9 cache: tuple of four tensors
10 lastkcache: (B, 1, D), last token k in current chunk (for RNN-style update)
11 lastvcache: (B, 1, D), last token v in current chunk (for RNN-style update)
12 kcache: (B, A, C, P) or (B, num_head, num_chunk, head_dim): cached k at each chunk’s end
13 vcache: (B, A, C, P) or (B, num_head, num_chunk, head_dim): cached v at each chunk’s end
14 # Note: KVCache is stored per chunk, not per token.
15 Returns:
16 out: (B, 1, D)
17 """
18 lastkcache, lastvcache, kcache, vcache = cache
19
20 # Intra-chunk RNN: Recurrent update within the current chunk
21 intra_k = g * lastkcache + (1.0 - g) * k
22 intra_v = g * lastvcache + (1.0 - g) * v
23 lastkcache.copy_(intra_k)
24 lastvcache.copy_(intra_v)
25
26 # Rearrange input for attention
27 intra_k = rearrange(intra_k, "b t (a p) -> b a t p", a=num_head)
28 intra_v = rearrange(intra_v, "b t (a p) -> b a t p", a=num_head)
29 q = rearrange(q, "b t (a p) -> b a t p", a=num_head)
30
31 # Apply RoPE
32 q, intra_k = apply_rope(q, intra_k)
33
34 # Update the current chunk’s kvcache
35 kcache[:, :, chunk_start:chunk_start + 1] = intra_k
36 vcache[:, :, chunk_start:chunk_start + 1] = intra_v
37
38 # Inter-chunk attention: standard attention over cached chunk representations
39 attn_out = scaled_dot_product_attention(
40 q, kcache[:, :, :chunk_start + 1], vcache[:, :, :chunk_start + 1], is_causal=False
41 )
42 out = rearrange(out, "b a t p -> b t (a p)", a=num_head)
43 return z * out
Listing 2: Pseudo code for the generation mode of RAT. We simply use the flash attention (Pytorch’s one). Note that KVCache of RAT is reduced from $T$ to $C$ compared to the attention module.
A.2 Experiments
Dataset
In our preliminary studies, we use the PG19 dataset [46], a long-form English book corpus with inherently long contexts. For the 1.3B model experiments, we adopt the FineWeb-Edu dataset [20], using its 100B-token randomly sampled version downloaded from the HuggingFace repository. To match the pretraining context length, we concatenate documents using a separator token. Note that web samples are usually very short, compared to the book dataset. For downstream evaluation, we consider a suite of classical commonsense reasoning benchmarks from the Eleuther AI evaluation harness [22], including PIQA [47], ARC-C [48], and HellaSwag [49]. In the LLaMA2 tokenizer, we observe that the inputs of these tasks typically contain fewer than 300 tokens. For the LongBench evaluation, in LABEL:tab:longbench we provide a mapping of task numbering to names and annotate their input lengths for each task, offering a rough indication of task difficulty with respect to input length. For SFT-based tasks, we have elaborated sequence lengths in the main text. As LongBench and SFT tasks often involve very long inputs, we apply truncation in the middle to preserve information at both the beginning and the end.
Table 5: We provide a mapping from task numbers to task names. NQA (NarrativeQA), MQA (MultiFieldQA-en), HQA (HotpotQA), WQA (2WikiMultihopQA), MSQ (Musique), GR (GovReport), MN (MultiNews), TQA (TriviaQA), and RBP (RepoBench-P). We also report the average input length and the 95th percentile input length of each task, measured in tokens using the LLaMA2 tokenizer.
Task Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
1-1 1-2 1-3 2-1 2-2 2-3 3-1 3-2 3-3 4-1 4-2 4-3 6-1 6-2
Name NQA Qasper MQA HQA WQA MSQ GR QMSum MN TREC TQA SAMSum LCC RBP
Avg. 36037 5780 8115 15329 8483 18555 12280 15980 3156 7792 14098 11168 4307 14818
95th pctl. 77966 10164 14994 19755 16939 20066 25721 29069 7135 12532 25006 19823 10401 31937
200M model
We start with a 200M-parameter model in our preliminary study, with a model dimension of 1024, 12 transformer layers, and head dimension of 64. The rotary position embedding (RoPE) base is set to 10,000. We use the GPT2 tokenizer. Following Mohtashami and Jaggi [45], we repeat the PG19 training split five times to reach a total of 15B training tokens. The learning rate is scheduled using cosine annealing, starting at $6.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with a warm-up ratio of 10%. We use the AdamW optimizer with a weight decay of 0.1 and $\beta=(0.9,0.98)$. Gradient clipping is applied with a threshold of 1.0, and the global batch size is set to 1M tokens. We explore training with three different context lengths: 8K, 16K, and 32K. Training these models on 4 H100 GPUs takes approximately 5 to 14 hours. In particular, the attention model requires up to 14 hours when the sequence length is $T=32768$, whereas the RAT(L=16) model completes training in about 7 hours under the same setting.
1.3B model
The 1.3B-parameter model uses a model dimension of 2048, 24 transformer layers, and a head dimension of 128, equipped with RMSNorm [50] and without bias. The RoPE base is also set to 10,000. The model parameters are initialized using a Gaussian distribution with a standard deviation of 0.02. We adopt the LLaMA2 tokenizer in the following studies. For pretraining, we use a cosine-annealed learning rate schedule starting at $8.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with 5% warmup. The global batch size is set to 2M tokens. We primarily train with a context window of 4K tokens and include additional experiments with a 16K context window, for which the RoPE base is increased to 500,000. Each model is trained on 16 H100 GPUs, requiring approximately 2 to 3 days to complete.
For LongBench evaluation, we follow the default prompts with greedy decoding for all tasks except summarization. For summarization, we apply a repetition penalty of 1.2 to address the common issue in pretrained-only models of generating repetitive outputs under instructional prompts. For SFT tasks, we train the models on the official training splits with an answer-only loss and evaluate them on the corresponding test sets. Although we explored different hyperparameters during the early experimentation, we observed that the relative trends remained largely stable. Thus, we fix the learning rate and batch size to $(1.0\times 10^{-5},128)$ for large datasets, and $(1.0\times 10^{-5},32)$ for the smaller QMSum [24] task, following common practice for 1B-scale models. The weight decay is set as 0.01, and all other hyperparameters follow the pretraining setup. We sweep over the number of epochs and report the best result for each dataset and architecture. In practice, QA tasks typically converge in 1–2 epochs, while summarization benefits from slightly longer training.
Appendix B Efficiency
B.1 Latency
To supplement Fig. 3 in main text, we put the concrete latency number of a single layer in Table 6, Table 7, and Table 8.
Table 6: Single layer (including input and output projections) training time across different sequence lengths. The latency (ms) is tested on 200K tokens.
Model 4096 8192 16384 32768 65536 131072 262144 RNN 77.00 78.42 77.50 83.50 115.83 130.62 195.30 RAT(L=128) 150.19 150.76 156.65 154.52 159.94 165.07 206.46 RAT(L=64) 146.25 148.46 151.10 155.83 165.58 177.78 227.76 RAT(L=16) 150.08 153.63 158.98 173.89 202.00 249.79 378.11 Attention 122.06 176.44 277.61 474.90 877.14 1713.82 3417.48 Attention/RAT(L=16) 0.81$\times$ 1.15$\times$ 1.75$\times$ 2.73$\times$ 4.34$\times$ 6.86$\times$ 9.04$\times$
Table 7: Single layer (including input and output projections) prefilling time across different sequences lengths. The latency (ms) is tested on 200K tokens.
Model 4096 8192 16384 32768 65536 131072 262144 RNN 24.93 24.75 25.32 27.09 30.54 39.40 56.16 RAT(L=128) 44.74 44.78 45.09 45.50 46.67 48.62 51.75 RAT(L=64) 43.46 44.27 44.27 44.83 47.08 50.83 57.38 RAT(L=16) 44.03 44.90 45.53 51.08 57.31 71.18 99.53 Attention 36.93 52.71 80.21 135.68 245.14 494.62 997.50 Attention/RAT(L=16) 0.84$\times$ 1.17$\times$ 1.76$\times$ 2.66$\times$ 4.28$\times$ 6.95$\times$ 10.02$\times$
Table 8: Single layer (including input and output projections) generation time at the specified position. The latency (ms) is tested on generating batches of tokens with $B=64$, $B=512$, and $B=1024$.
Model 4096 8192 16384 32768 65536 131072 262144 $B=64$ RNN 0.36 0.33 0.34 0.33 0.33 0.33 0.33 RAT(L=128) 0.62 0.60 0.61 0.63 0.66 0.66 0.68 RAT(L=64) 0.67 0.64 0.66 0.70 0.73 0.74 1.03 RAT(L=16) 0.62 0.65 0.64 0.68 1.04 1.82 3.42 Attention 1.15 1.73 3.33 6.66 12.94 26.8 OOM $B=512$ RNN 0.35 0.34 0.33 0.33 0.34 0.34 0.33 RAT(L=128) 0.63 0.64 0.66 0.75 1.17 2.44 3.46 RAT(L=64) 0.70 0.68 0.74 1.16 2.41 3.44 6.69 RAT(L=16) 0.75 1.22 2.45 3.48 6.73 13.09 25.76 Attention 6.56 12.89 25.84 OOM OOM OOM OOM $B=1024$ RNN 0.36 0.36 0.35 0.35 0.35 0.36 0.35 RAT(L=128) 0.73 0.75 0.95 1.35 2.54 4.64 6.98 RAT(L=64) 0.74 0.94 1.35 2.48 4.67 7.02 13.45 RAT(L=16) 1.38 2.56 4.63 7.04 13.51 26.43 OOM Attention 13.04 25.70 OOM OOM OOM OOM OOM
B.2 Maximum throughput
As RAT also reduces cache memory usage, we report the maximum throughput of the full 1.3B model across different sequence lengths in Fig. 5. For example, RAT(L=16) achieves an approximately 10.2×10.2$\times$10.2 × higher throughput than the baseline attention model at $T=4096$, and improves further to 15.6×15.6$\times$15.6 × at $T=16384$. Similarly, RAT(L=16)-SWA reaches 4.5×4.5$\times$4.5 ×, 8.0×8.0$\times$8.0 ×, and 13.9×13.9$\times$13.9 × higher throughput over attention at $T=4096$, $8192$, and $16384$, respectively.
Figure 5: We measure the maximum throughput of the full 1.3B model for generating 1024 tokens under different prefilling lengths. For each total sequence length $T$, the prefilling length is set to $T-1024$. For example, $T=4096$ corresponds to a prefilling of 3072 tokens, while $T=8192$ and $T=16384$ correspond to 7168 and 15360 tokens, respectively. When the sequence length increases, the maximum throughput ratio between RAT and attention rises from 10.2×10.2$\times$10.2 × to 15.6×15.6$\times$15.6 ×, highlighting the strong efficiency advantage of RAT in long-context generation.
Appendix C Accuracy
C.1 Reversed design of RAT
Instead of employing recurrence within each chunk, we also explored a reversed design that applies attention inside the chunk, followed by recurrence to capture long-range dependencies. For this reversed design, standard RoPE is sufficient. However, with respect to parameter allocation, we found it crucial to retain the key vector in the attention module as a full tensor, as sharing it together with the query vector leads to a collapse into single-head attention. Thus, our final reversed design shares the query vector across attention heads, employs low-rank matrices for the projection used for output gate, and preserves full projections for both the forget gate and the key vector. Results of the reversed design are put in Table 9. It can be observed that, under identical FLOP constraints, this design significantly underperforms RAT. For example, to match the FLOPs budget of $\mathcal{O}(T/64)$, it requires chunk sizes of 128, 256, and 512 for sequence lengths of 8192, 16384, and 32768, respectively. Most of the computational cost is spent within a small local window of 128 tokens, while excessive burden is placed on the inter-chunk recurrence, ultimately leading to high perplexity. Only when we reduce the number of chunks to 16 does the perplexity drop below 14.00. Therefore, while the reversed design also allows for interpolation between attention and RNN (reducing to attention when $L=T$, and to RNN when $L=1$), we choose to focus on RAT due to its more efficient utilization of FLOPs.
Table 9: Perplexity results of the reversed design, where attention is applied within chunks and RNN is used across chunks. Experiments are conducted on the 200M model trained on the PG19 book dataset. Under the same FLOPs, it can be observed that RAT significantly outperforms its reversed counterpart.
Method FLOPs $\mathbf{T=8192}$ $\mathbf{T=16384}$ $\mathbf{T=32768}$ RAT(L=128) $\mathcal{O}(T/128)$ 13.57 13.44 13.66 RAT(L=64) $\mathcal{O}(T/64)$ 13.42 13.34 13.39 RAT-Reversed(C=128) $\mathcal{O}(T/128)$ 14.53 14.35 14.37 RAT-Reversed(C=64) $\mathcal{O}(T/64)$ 14.21 14.06 14.02 RAT-Reversed(C=16) $\mathcal{O}(T/16)$ 13.69 13.48 13.50 Attention $\mathcal{O}(T)$ 13.26 13.10 13.19
Table 10: Pretraining and downstream evaluation results of 1.3B models using either RoPE or NoPE positional encodings. RoPE is used in the main text, while NoPE is trained for investigating length extrapolation. Notably, NoPE achieves reasonable performance even at the 1B model scale with only 100B training tokens.
Method PE Pretrain HellaSwag LAMBADA PIQA PPL acc_norm acc acc_norm Attention-SWA RoPE 7.61 57.1 48.0 72.7 RAT(L=16)-SWA RoPE 7.57 57.9 49.3 73.9 Attention-SWA NoPE 7.69 56.4 47.3 72.7 RAT(L=16)-SWA NoPE 7.63 56.8 47.8 73.7
C.2 Length generalization
Figure 6: Evaluation at different test lengths for pretrained models trained with a 4K context window. RAT(L=16)-SWA with NoPE achieves the best overall performance, exhibiting strong generalization up to $T=16384$ while maintaining low loss within the training context.
Because of the use of softmax-based attention and RoPE at the inter-chunk level, it is reasonable to expect that RAT may also face challenges in length generalization, where the model is trained on short sequences but evaluated on much longer ones. To study this, we consider SWA variants using either RoPE or NoPE for attention and RAT, following recent practices that interleave attention with local attention modules and apply positional encodings only in the local attention. This design has been adopted in practice and has shown promising results [51].
As shown in Fig. 6, the RNN performs very steadily as the test sequence length increases. However, within the training context, it has the highest loss among all models. With RoPE, both Attention-SWA and RAT(L=16)-SWA experience a sharp increase in loss when the sequence length goes beyond 6000. The increase is more severe in the attention model than in the RAT variant. This is likely because RAT reduces the effective attention span by attending only to inter-chunk positions, which may offer better robustness to extrapolation.
We also pretrain NoPE versions, where no positional encodings are used in attention or RAT layers. Interestingly, these NoPE models show reasonable performance, as reported in Table 10. While their pretraining perplexities are higher than those of the RoPE variants, the gap remains small, especially considering the 1.3B model size and 100B-token training budget. To further verify this, we evaluate the models on several commonsense reasoning tasks and obtain promising results. We expect that NoPE will perform even better with larger models and longer training. In terms of length extrapolation, using NoPE significantly improves the robustness for both Attention-SWA and RAT(L=16)-SWA. Among them, RAT(L=16)-SWA with NoPE shows the most stable performance, maintaining low loss even at a sequence length of 16,384.
In conclusion, we find that RAT can outperform the attention module in length generalization, both with RoPE and with NoPE. Still, the problem is not fully solved, as none of the models reach the stability level of the simple RNN. While many techniques have been proposed to improve length extrapolation in attention models, they should also be considered for RAT, as RAT essentially redirects attention from every position to inter-chunk locations. For example, RoPE extrapolation in the attention module can be improved using methods from Peng et al. [52], bloc97 [53], and NoPE extrapolation has been studied in Wang et al. [54], which points out that NoPE still has a context length limit, though it performs better than RoPE. Their work links the failure to shifts in attention distributions and proposes tuning the temperature of attention heads to improve extrapolation. We believe such methods could also be applied to RAT, and leave further exploration of these techniques to future work.
C.3 16K context window
Table 11: Performance of 1.3B models with a 16K context window, including pretraining perplexity, commonsense reasoning accuracy (short-context), and SFT results on long-context tasks. For SFT, we focus on tasks that still have long context even with the 16K context window, such as QMSum and NarrativeQA3. For NarrativeQA3, we adopt the passage-only setting from the original paper instead of the summary+passage setup used in NarrativeQA2. Although passages are still truncated at 16K, most information are preserved, making passage-only a more meaningful and challenging evaluation setting.
Model Throughput Pretrain Evaluation SFT
token/sec Val. ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg. NarrativeQA3 QMSum
RNN 130K 8.09 37.4 65.7 54.2 37.6 72.6 53.4 53.5 15.4 22.8
RAT(L=16) 10376 7.52 36.5 64.8 55.8 45.6 71.5 53.5 54.6 21.6 24.2
Attention 667 7.44 36.8 67.5 57.0 43.9 73.2 57.1 55.9 23.3 23.8
Attention-SWA 1347 7.45 36.9 65.8 56.6 47.6 72.5 55.6 55.8 20.3 23.3
RAT(L=16)-SWA 9244 7.42 38.5 67.5 57.6 47.7 73.3 57.0 56.9 22.6 24.6
We also pretrain the models with a 16K context length. However, we observe that directly pretraining with such a long context window may lead to issues, as most web data is relatively short and needs to be concatenated by the separate token to fill the window. Due to the lack of naturally long-context data, the model may fail to learn true long-context capabilities. We believe this setting would benefit from more careful strategies, such as continued pretraining and balanced long-context data sources. Nonetheless, due to resource limitations, we directly pretrain the models on a 16K context length in this work, and hope that the results can reflect the ability of different architectures to handle longer contexts to some extent.
Due to the short web data used during pretraining, we consider evaluating the models on short-context tasks, or applying SFT on long-context tasks that contain real long sequences. As shown in Table 11, overall, the performance of RAT(L=16) stays in between of RNN and attention, and RAT(L=16)-SWA performs the best. It achieves a 0.02 pretraining lower perplexity than attention, and improves by about 1 point on average in commonsense reasoning and 0.8 points on QMSum. However, it underperforms attention around 0.7 points in F1 score on the NarrativeQA3 task, unlike in the 4K context length setting. On the other hand, the maximum throughput improves significantly: from 4.5$\sim$ 4.5$\times$ 4.5 × compared to attention at the 4K context window to 13.9$\sim$ 13.9$\times$ 13.9 × at 16K. This suggests that, to pursue higher accuracy performance, we may consider increasing the local attention window size (e.g., from 1024 to 4096) under the 16K context length setting.
C.4 Needle-in-Haystack (retrieval ability)
To evaluate retrieval capabilities—an area where attention-based architectures are known to excel, and where RNNs typically fall short—we conduct experiments on the Needle-in-Haystack tasks. In these tasks, a “needle” (e.g., a magic number or UUID) is embedded within irrelevant passages, noisy text, or mixed with multiple key-value pairs. Each key-value pair is presented as a short identifier (the key) followed by its associated value, and the model is prompted to retrieve the correct value given a specific key.
We adopt the RULER benchmark [55] and evaluate a range of Needle-in-Haystack task configurations. Specifically, task single_1 involves retrieving a specific number from a background filled with repeated noise. In single_2, the background is replaced with natural stories. single_3 increases the difficulty by requiring retrieval of a long and complex UUID. The multikey tasks are more challenging than the single-key ones, as they introduce multiple such key-value pairs into the context. In particular, multikey_2 and multikey_3 consist almost entirely of densely packed key-value pairs, among which only a single key is queried at the end. This makes the task especially challenging, as the model must retrieve the correct item from a highly cluttered input full of distractors. And multikey_3 further incorporates the complex UUID format. In the multiquery and multivalue settings, the model is required to resolve multiple retrieval targets—either by answering several distinct queries, or by retrieving all values associated with a single key. For detailed definitions of each task, we refer the reader to the benchmark documentation.
During evaluation, we observed that models may fail to interpret certain prompts correctly. For instance, prompts like “A special magic number is hidden within the following text. Make sure to memorize it. I will quiz you about the number afterwards.” can cause failures even in attention-based models, especially on the harder multikey tasks. To mitigate this, we apply a light, one-round supervised fine-tuning stage before evaluation to adapt the models to the instruction patterns. We generate 1,000 synthetic training samples for each of the 8 tasks, resulting in a total of 8,000 examples disjoint from the validation sets. The models are trained on this dataset for one round and then are evaluated directly on all 8 tasks. This procedure yields a fairer and more stable comparison. As shown in Table 12, RNNs perform reasonably well on the simpler tasks (single_1, single_2), but their performance drops to near zero on the harder ones. Attention-based models perform consistently well, benefiting from their ability to access all previous tokens directly. RAT(L=16) achieves results close to full attention, especially in numeric tasks. However, it struggles on UUID tasks due to their complexity and the use of exact match scoring (single_3, multikey_3). Meanwhile, RAT(L=64) falls between RNNs and RAT(L=16), as expected given its partial access to long-range context.
These results are consistent with the underlying architecture designs: attention provides full direct access to all past tokens, RNNs compress all past information into a hidden state, while RAT compresses part of the history but also retains direct access at the chunk level. As a result, RAT naturally exhibits retrieval capabilities that lie between those of attention and RNNs.
Table 12: Accuracy on the Needle-in-Haystack tasks with different configurations from the RULER benchmark [55], used to evaluate retrieval capabilities. Evaluation is conducted with a 4K context window, using exact match as the metric.
Model single_1 single_2 single_3 multikey_1 multikey_2 multikey_3 multiquery multivalue
RNN 99.4 99.8 0.0 2.2 0.0 0.0 5.8 5.7
RAT(L=64) 100.0 99.8 55.6 96.4 62.2 1.0 86.3 89.1
RAT(L=16) 99.6 100.0 94.6 99.6 99.6 82.6 91.2 94.8
Attention 100.0 100.0 100.0 99.6 99.6 95.2 98.6 99.1
Appendix D Broader Impacts
Enhancing the efficiency of Large Language Models (LLMs) can significantly reduce computational resources and energy consumption, benefiting the environment and democratizing access to advanced AI technologies. However, increased efficiency could also lead to greater dissemination of disinformation and the creation of deepfakes, posing risks to public trust and security and potentially reinforcing existing biases that impact specific groups unfairly. This research aims to promote the responsible development and deployment of LLMs, maximizing societal benefits while acknowledging potential harms.
Appendix E License information

### v3

@@
-RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling
+RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy via Chunk-based Sequence Modeling
 Abstract
-Transformers have become the cornerstone of modern large-scale language models; however, their dependence on softmax attention poses a major computational bottleneck, particularly in long-context settings.
-In this work, rather than following prevalent approaches such as linear attention (or SSMs) and local attention, we introduce an intermediate design called RAT between recurrence and attention mechanisms.
-It partitions the input into chunks, applies a simple linear recurrence within each chunk to capture local dependencies, and then performs softmax attention across chunks to model long-range interactions.
-By adjusting the size of the chunk, RAT enables flexible trade-offs, combining the strengths of RNN and attention.
-Empirically, with a chunk size of 16, the RAT layer achieves a 7×7$\times$7 × improvement in training speed with 100K token sequences and 9×9$\times$9 × in generation at 4K sequence length, while maintaining similar or sometimes even better accuracy compared to standard attention.
+Transformers have become the cornerstone of modern large-scale language models, but their reliance on softmax attention poses a computational bottleneck at both training and inference.
+Recurrent models offer high efficiency, but compressing the full sequence into a fixed-size and holistic representation can suffer from memory degradation in long contexts and limit fine-grained retrieval.
+To address this, we propose RAT, an intermediate design that bridges the efficiency of RNNs and capacity of attention.
+RAT partitions the input into chunks, applies recurrence within each chunk for local dependencies, and softmax-based attention across chunks for long-range interactions.
+This design mitigates memory degradation and enables direct access to distant tokens, while retaining computational efficiency.
+Empirically, with a chunk size of 16, the RAT block achieves a $7\times$ improvement in training speed for 100K sequence length and $9\times$ in generation at the 4K position, while maintaining similar performance compared to standard attention.
 We demonstrate this by training 1.3B parameter models from scratch and performing large-scale evaluations, including short- and long-context benchmarks, as well as supervised fine-tuning (SFT).
 We further propose a hybrid architecture that interleaves RAT with local attention.
-By combining efficient long-range modeling with strong local interactions, this hybrid design not only improves inference speed and reduces cache memory usage compared to attention, but also consistently enhances performance, for example, achieving an average 1 point gain in commonsense reasoning tasks, up to 4 points on code tasks, and a 1 point Rouge-L increase in a summarization SFT task.
+By combining efficient long-range modeling with strong local interactions, this hybrid design not only improves inference speed and reduces cache memory usage, but also consistently enhances performance and shows the overall best results.
 Code is available at https://github.com/CLAIRE-Labo/RAT.
 1 Introduction
-Language modeling has long been dominated by Transformer-based architectures because of their remarkable performance across a wide range of tasks.
-However, their reliance on full self-attention leads to quadratic complexity with respect to sequence length, which limits scalability in long-context processing.
-This has motivated many efforts to design more efficient attention mechanisms.
-Some approaches focus on local attention [1, 2, 3], and use sparse attention or recurrence to capture long-range information.
-Another line of work investigates the duality between linear RNNs and attention [4, 5, 6, 7], by enforcing strict linearity constraints and introducing the notion of state expansion, which represents recurrent states as matrices rather than vectors, and shifting from per-dimension to per-head gating.
-In contrast to them, we propose an alternative architecture.
-We argue that overusing attention in short-range contexts underutilizes its strengths and that such local processing can instead be handled more efficiently by lightweight structures such as recurrences.
-Based on this insight, we explore a design that lies between RNNs and attention.
-This design also avoids strict linearity assumptions on either side and adopts the classical recurrence form with per-dimension gating and no state expansion.
-Figure 1: Comparison of three architectures: The circles represent the token representations and the arrows indicate the information flow for the 5th token.
-(a) Attention: Each token retains its representation and has full access to all previous tokens.
-(b) RNN: The information is progressively compressed along the sequence, and the current token only accesses the last hidden state.
-(c) RAT(L=2): Intra-chunk recurrence compresses local information, while inter-chunk attention allows direct access to previous chunks.
-Note that (a) and (b) can be viewed as special cases of RAT with $L=1$ and $L=T$, respectively.
-Figure 2: RAT structure. The symbol $\sigma(\cdot)$ further applies a sigmoid function to the gates. Note that we do not use extra convolution or normalization in our design.
-We propose a RAT layer, which is a simple but effective temporal mixing mechanism.
-We divide long sequences into chunks: within each chunk, we apply recurrence for local modeling, and use softmax-based attention across chunks to access distant information directly (see Fig.
-1).
-Recurrence efficiently captures short-range dependencies without suffering from the memory degradation issues common in long sequences, while attention over chunk-level representations enables long-range retrieval.
-Adjusting the size of the fragment $L$, RAT interpolates between attention (when $L=1$) and RNN (when $L=T$).
-To make training more efficient, we adopt a lightweight gated linear recurrence [8].
-We further explore a hybrid design that interleaves RAT with sliding-window attention layers [1], allowing the two mechanisms to complement each other while remaining efficient.
-In terms of efficiency, our implementation relies solely on PyTorch high-level operators without any custom CUDA kernels, yet already achieves a substantial speed-up.
-With a chunk size of $L=16$, the single layer latency to generate the token at position 4096 is up to 9×9$\times$9 × faster than that of vanilla attention.
-And as shown in Table 1, the maximum throughput of the full model improves by up to 10×10$\times$10 × over attention, and interleaving with local attention can also yield a 4×4$\times$4 × gain.
-We pretrained models with different chunk sizes at the 1.3B scale using a 4K context length, and found their perplexity (PPL) falls between that of RNNs and attention.
-We then evaluated them in three settings: (1) zero-shot commonsense reasoning with short contexts, (2) long-context understanding on LongBench [9], and (3) SFT on two question answering (QA) and two summarization tasks.
-RAT with $L=16$ performs comparably to attention on most benchmarks and even surpasses it on some LongBench tasks.
-When interleaved with local attention, it achieves the strongest overall results across all variants while maintaining high efficiency.
+Language modeling has long been dominated by Transformer-based architectures due to their strong performance across a wide range of tasks.
+However, their reliance on full self-attention [attention] results in quadratic time and memory complexity with respect to sequence length, which limits scalability in long-context processing.
+This limitation has motivated a wave of recent efforts to revisit recurrent models or propose novel linear recurrent models such as state space models, linear attention methods [mamba, mamba2, linear_attention, gateddeltanet, xlstm, griffin, lru].
+By comparing these architectures, we observe a key difference between recurrent models and self-attention.
+Recurrent approaches compress the full sequence history into fixed-size and holistic representations, which can lead to degraded memory when modeling long sequences and limit precise information retrieval.
+In contrast, self-attention retains full-token access and thus do not suffer from the two problems, but at the cost of heavy computation.
+This motivates us to explore an intermediate design that partially compresses the sequence while still maintaining global access.
+We propose a RAT layer, a simple yet effective temporal mixing method with a chunk-based design.
+It divides long sequences into chunks, applying recurrence within each chunk for local modeling and softmax-based attention across chunks for direct access to distant information (see 1(a)).
+Recurrence efficiently captures short-range dependencies while avoiding the memory degradation common in long sequences, whereas attention over chunk-level representations enables long-range retrieval.
+By adjusting the chunk size $L$, RAT interpolates between attention (when $L=1$) and RNN (when $L=T$).
+(a) Comparison of three structures: Circles denote token representations and arrows show the information flow for the 5th token.
+(1) Attention: Each token keeps its own representation and has full access to all previous tokens.
+(2) RNN: Information is progressively compressed along the sequence, and the current token accesses only the last hidden state.
+(3) RAT(L=2): Intra-chunk recurrence compresses local information, while inter-chunk attention enables direct access to previous chunks.
+(1) and (2) can be viewed as RAT with $L=1$ and $L=T$, respectively.
+(b) RAT structure. The symbol $\sigma(\cdot)$ further applies a sigmoid function to the gates.
+To make RAT scalable and efficient, we further explore its positional encoding, parameter allocation, and efficient implementations that specifically address the causal masking problem in this chunk-based design.
+Moreover, we investigate a hybrid architecture that interleaves RAT with sliding-window attention [longformer, gau, transformer-xl, block-recurrent], which focuses computation on local windows.
+As RAT suggests that overusing attention in local contexts underutilizes its strengths, and such dependencies can be handled more efficiently via lightweight recurrence, the two approaches can be complementary.
+We demonstrate that RAT is both efficient and performant.
+For instance, with a chunk size of $L=16$, the latency of a single temporal mixing block at position 4096 is up to $9\times$ lower than that of attention.
+As shown in Table 1, the full model achieves up to $10\times$ higher maximum throughput, and interleaving with local attention yields about $4\times$ improvement.
+We pretrain models at 1.3B scale and compare their results on 1) zero-shot commonsense reasoning with short contexts, (2) long-context understanding on LongBench [longbench], and (3) supervised fine-tuning on long-context tasks.
+RAT with $L=16$ performs on par with full attention in most benchmarks and even outperforms it on several LongBench tasks.
+When interleaved with local attention, it achieves the best overall results across all variants while maintaining high efficiency.
 Our main contributions are as follows:
   1. 1.
-    We propose the RAT layer as an intermediate design between RNN and attention for temporal mixing.
-We explore its detailed design and efficient implementation across three key scenarios, and observe significant speed.
+    We propose the RAT layer, a novel intermediate architecture that bridges the efficiency of recurrence and the capacity of attention.
+It compresses only local context while preserving global access, enabling direct retrieval and avoiding the memory degradation caused by full sequence compression in long contexts.
   2. 2.
-    We pretrain models at the 1.3B scale and evaluate them on 6 short-context, 14 long-context, and 4 SFT tasks, showing that RAT(L=16) performs competitively with attention.
+    RAT is simple, scalable, and efficient.
+It requires no custom CUDA kernels and is naturally compatible with existing multi-dimension parallelism schemes.
+We also introduce a hybrid variant by interleaving RAT with sliding-window attention, enabling efficient long-range modeling with strong local interactions.
   3. 3.
-    We further investigate a hybrid variant that interleaves RAT with sliding-window attention, which yields the best overall performance—for example, +1 accuracy on commonsense reasoning tasks, +4 on code completion, +4 on a challenging QA SFT task, and +1 on a challenging summarization task—compared to full attention or its local variant.
+    We validate RAT through extensive experiments at the 1.3B scale across diverse tasks, including 7 short-context reasoning benchmarks, 11 long-context tasks, 4 supervised fine-tuning objectives, and 9 retrieval-heavy synthetic evaluations.
+RAT with L=16L{=}16 shows comparable performance to full attention with $9\times$ faster single-layer decoding.
+Its hybrid variant with local attention yields the best overall results, along with a $3-4\times$ maximum throughput boost—for instance, +1 accuracy on commonsense reasoning, +4 on code completion, +4 on a challenging QA task, and +1 on a difficult summarization task.
 Table 1: Representative results for 1.3B models across pretraining, direct evaluation, and SFT.
@@
 Rouge-L Avg.
-EditSum F1 Rouge-L Attention 3052 7.61 56.7 18.2 19.5 23.9 61.3 23.4 RAT(L=16) 31170 7.67 55.6 19.6 20.2 17.4 60.8 23.3 Attention-SWA 4605 7.61 56.4 17.4 19.4 21.7 63.3 23.4 RAT(L=16)-SWA 13582 7.57 57.8 18.8 19.5 28.2 63.2 24.6
-2 Method
-2.1 Attention in temporal mixing
+EditSum F1 Rouge-L Attention 3052 7.61 56.9 18.2 19.5 23.9 61.3 23.4 RAT(L=16) 31170 7.67 56.7 19.6 20.2 17.4 60.8 23.3 Attention-SWA 4605 7.61 57.1 17.4 19.4 21.7 63.3 23.4 RAT(L=16)-SWA 13582 7.57 58.0 18.8 19.5 28.2 63.2 24.6
+2 Overall architecture
+To motivate our design, we first examine how attention and recurrence compress contextual information at the token level, along with the equations that will be reused later.
+We then introduce RAT as an intermediate mechanism that inherits the respective advantages of both.
+2.1 Attention: Full-Token Access
 In softmax attention mechanisms, each token has access to all preceding tokens through a learned token-specific weighted aggregation:
 ${\bm{y}}_{t}=f({\bm{q}}_{t}{\bm{K}}^{\top}_{:}){\bm{V}}_{:}$ (1)
-$f(\cdot)$ denotes causal masking and the softmax function, and $f({\bm{q}}_{t}{\bm{K}}^{\top}_{:})$ are the attention weights used to aggregate ${\bm{V}}_{:}$.
-We use Python-style indexing in ${\bm{K}}^{\top}_{:},{\bm{V}}_{:}$ to highlight that each query is concerned with all keys and values.
-Although attention is widely considered a crucial component in state-of-the-art architectures for sequence modeling, its full-token access incurs a computational complexity of $\mathcal{O}(T^{2})$, where $T$ is the length of the sequence.
-This quadratic growth in FLOPs can lead to substantial inefficiency during training and inference.
-We view attention as a slow but lossless module in the mixing of information between tokens and aims to combine it with a fast but lossy component.
-2.2 Recurrence in temporal mixing
-RNNs exploit sequential dependencies in the data [10, 11].
-As the model processes the sequence, it maintains a summary of past information in its hidden state, allowing each step to be computed solely from the previous state and the current input.
-This enables efficient inference and low FLOPs, in contrast to the full-sequence access required by attention.
-To further improve training efficiency, recent work has explored linear RNN variants [12, 13].
-For simplicity, we adopt the design proposed in Feng et al.
-[8], which is similar to that in Smith et al.
-[14].
-Using diagonal matrices and removing nonlinearities, the recurrence reduces to an EMA-like gating behavior, which can be efficiently implemented with the parallel scan algorithm [14]:
+$f(\cdot)$ denotes causal masking and the softmax function, and $f({\bm{q}}_{t}{\bm{K}}^{\top}_{:})$ represents the attention weights used to aggregate $\bm{V}$$_{:}$.
+We adopt Python-style indexing in ${\bm{K}}^{\top}_{:},{\bm{V}}_{:}$ to emphasize that each query attends to all keys and values.
+This full-token access makes attention a dominant architecture in sequence modeling, but it suffers from high computational cost in both training and inference.
+2.2 Recurrence: Full-Sequence Compression
+Recurrent models maintain a summary of past information in a fixed-size representation, allowing each step to depend only on the previous state and the current input [gru, lstm].
+To initiate the design of RAT, we adopt a simple and fast linear recurrence [minrnn, simple_linear_rnn, s5], which uses diagonal matrices and no nonlinearities to simplify the classic recurrence into an EMA-like gating mechanism:
 $\tilde{{\bm{v}}}_{t}$ $={\bm{g}}_{t}\odot\tilde{{\bm{v}}}_{t-1}+(1-{\bm{g}}_{t})\odot{\bm{v}}_{t},$ (2)
 ${\bm{y}}_{t}$ $={\bm{z}}_{t}\odot\tilde{{\bm{v}}}_{t},$
-where ${\bm{g}}_{t}$ and ${\bm{z}}_{t}\in\mathbb{R}^{D}$ denote the forget and output gates, respectively, computed via linear projections of the input followed by a sigmoid activation.
+where ${\bm{g}}_{t}$ and ${\bm{z}}_{t}\in\mathbb{R}^{D}$ denote the per-dimension forget and output gates, respectively, computed via linear projections of the input followed by a sigmoid activation.
 Here, $D$ refers to the dimension of the model.
-We find this simple design sufficient to achieve strong performance for short sequences.
-However, compressing the entire history to the current state, even with more complex recurrences, can lead to memory degradation over time [15].
-As a result, RNN may struggle to retain information from distant time steps, which poses challenges for long-context tasks.
-Thus, in this paper, we place RNN as the fast yet lossy token mixing module.
-2.3 RAT: a design between RNN and attention
-Motivated by the complementary strengths of RNN and attention, we propose an intermediate design that combines their advantages: the speed of RNN and the global access of attention.
-Since RNN are effective for short sequences, while attention excels at modeling long-range dependencies, we reinterpret the input as a sequence of shorter chunks of sequences.
-A recurrent module is applied within each chunk to capture local dependencies, followed by cross-chunk attention to model global interactions efficiently, as shown in Fig.
-1.
-We divide a sequence of length $T$ into $C$ chunks of length $L$, such that $T=C\cdot L$.
+This can be efficiently implemented using the parallel scan algorithm [s5].
+We take this minimal choice to highlight the core idea, but the recurrence in RAT is not limited to this certain form and can be extended to more advanced variants such as 2D recurrence or nonlinear RNNs.
+We leave this as future work.
+The state-based formulation of recurrence yields low computational cost and efficient inference, and performs well on short sequences.
+However, compressing the entire sequence history into a fixed-size and holistic state, even with more expressive designs, can still suffer from memory degradation when the sequence length grows [rnn_difficulty], and limits precise information retrieval, particularly in noisy contexts.
+2.3 RAT: Chunk-Based Intermediate Design
+To bridge the strong performance of attention enabled by full-token access with the efficiency of RNNs derived from full-sequence compression, we propose an intermediate design by reinterpreting the input as a sequence of shorter chunks.
+A recurrent module is applied within each chunk to model local dependencies, followed by cross-chunk attention to enable global interactions, as illustrated in 1(a).
+This design mitigates the fixed-size representation limitation of RNNs and the inefficiency of attention.
+Technically, we divide a sequence of length $T$ into $C$ chunks of length $L$, such that $T=C\cdot L$.
 A token originally at position $t$ is re-indexed as $(c,l)$, where $c$ denotes the chunk index and $l$ the position within the chunk.
 Within each chunk, a forget gate ${\bm{g}}_{c,l}$ is used to recurrently aggregate the value ${\bm{v}}_{c,l}$ and key ${\bm{k}}_{c,l}$ vectors, yielding updated representations $\tilde{{\bm{v}}}_{c,l}$ and $\tilde{{\bm{k}}}_{c,l}$.
-For each query ${\bm{q}}_{c,l}$, we compute attention on the final key vectors $\tilde{{\bm{K}}}_{:,-1}$ from all preceding chunks, as well as the current chunk’s key $\tilde{{\bm{k}}}_{c,l}$.
+For each query ${\bm{q}}_{c,l}$, we compute attention on the chunk-level key and value vectors, including $\tilde{{\bm{K}}}_{:,-1}$ for all preceding chunks, and $\tilde{{\bm{k}}}_{c,l}$ for the current chunk.
 The causal masking function $f(\cdot)$ restricts attention to the chunks before it, followed by a softmax operation.
@@
 ${\bm{y}}_{c,l}={\bm{z}}_{c,l}\odot{\bm{y}}_{c,l}$
-The overall computation flow is illustrated in Fig.
-2.
-Benefiting from the short length of each chunk, the RNN is both effective and computationally efficient, since its hidden state has sufficient capacity to encode short-range context without discarding much information.
-The attention applied across chunks then enables access to distant positions without requiring compression.
-As a result, the overall design preserves accuracy comparable to that of full softmax attention, while reducing computational cost, as inter-chunk attention operates over shorter contexts, with attention length inversely proportional to the chunk size.
-We position this design as an intermediate architecture design between RNN and attention, aiming to combine their respective strengths, and refer to it as RAT.
-By adjusting the size of the chunk $L$, RAT interpolates between the RNN and the attention behaviors: when $L=T$, the model reduces to an RNN, since the entire sequence is treated as a single chunk; when $L=1$, it resembles the attention, since each token resides in its chunk.
-We also explored a reversed variant that applies attention within chunks and recurrence across chunks.
-Although this design also allows interpolation, it underperforms RAT with equal FLOPs, likely due to suboptimal allocation of computation to local attention.
-Further discussion is provided in Appendix C.
-In the following two paragraphs, we discuss the detailed design choices for RAT.
+The overall computation flow is illustrated in 1(b).
+Thanks to the short length of each chunk, the recurrent module efficiently captures local dependencies without significant information loss.
+Cross-chunk attention then enables global access without compressing distant information.
+This design preserves the accuracy of softmax-based attention while reducing computation, as inter-chunk attention operates over shorter sequences with $L$ as the FLOPs reduction ratio.
+We also experimented with a reversed variant by applying attention within chunks and recurrence across them, but found that standard RAT achieves better FLOPs utilization and overall performance (see Appendix C).
+Benefits of the chunk design
+We position this as an intermediate architecture between RNNs and attention, and refer to it as RAT.
+By adjusting the chunk size $L$, RAT interpolates between their behaviors: when $L=T$, it reduces to a pure RNN; when $L=1$, it resembles full attention.
+This allows us to pursue the trade-off between attention and RNN with a single-layer design, and offers greater flexibility for hybrid modeling by using different chunk sizes to emulate varying behaviors.
+From a mechanistic perspective, unlike either classic or advanced recurrence (state space or linear attention models) that rely on fixed-size and holistic representations, our chunk-based structure enables memory capacity to scale with sequence length while maintaining a fixed FLOPs reduction ratio.
+Its partial compression with direct access to prior chunks ensures its superior performance on retrieval-heavy tasks.
+3 Scalable and Efficient Modeling with RAT
+3.1 Design details
 Parameter allocation
-We aim to match the parameter budget of RNN and attention.
-In addition to the projection of the output, which introduces the parameters $D^{2}$ in all three mechanisms, both attention and RNN use additional parameters $3D^{2}$.
-These correspond to projections for the query, key, and value in attention, and for the value, the forget gate, and the output gate in RNN.
-To maintain the exact total parameter count for RAT, which combines both mechanisms, we adopt lightweight designs.
-Specifically, we consider using low-rank projections for the gates ${\bm{g}}$ and ${\bm{z}}$, or sharing query and key projections across heads.
+We aim to control the parameters of RAT at $4D^{2}$, given $D^{2}$ for the output projection, $3D^{2}$ for the query, key, and value projections in attention, and an additional $2D^{2}$ for the two gates in RNN.
+We explore lightweight alternatives by using low-rank projections for the gates $\bm{g}$} and $\bm{z}$}, or by sharing query and key projections across heads.
 Empirical results in Subsec.
-3.3 show that sharing query ${\bm{q}}$ and key ${\bm{k}}$ typically improves performance by enabling more efficient parameter allocation.
-Importantly, this design does not reduce to single-head attention because the forget gate operates at a per-dimension level and introduces different elements into the gated key $\tilde{{\bm{k}}}$ used for the inter-chunk attention.
+4.2 show that sharing the query $\bm{q}$} and key $\bm{k}$} slightly outperforms using low-rank gates.
+Notably, this design does not collapse into single-head attention, since the forget gate operates at the per-dimension level and produces distinct gated keys $\tilde{{\bm{k}}}$ for the inter-chunk attention.
 Positional encoding
-We examine how to encode positional information in RAT.
-Although RNN captures position through its sequential structure, the attention is usually based on explicit encodings such as RoPE [16].
-For RAT, we find that applying RoPE based on chunk indices rather than original token positions yields slightly better performance, and we use cross-chunk RoPE by default, as it may better align with inter-chunk attention.
-See Subsec.
-C.2 for a more detailed discussion, including exploration of the NoPE [17] encoding.
+We examine how to encode positional information in RAT, due to the presence of cross-chunk attention.
+Motivated by the fact that RNN captures positional information through its sequential structure, we find that applying positional encoding at the chunk level, rather than relying on original token positions, yields slightly better fidelity.
+This strategy also improves length generalization, as the number of positions requiring encoding (i.e., the number of chunks) is much smaller than the full sequence length.
+In our main experiments, we use RoPE [rope] based on chunk indices for inter-chunk attention.
+In the length generalization study (see Subsec.
+C.2), we further explore NoPE [nope], which yields the best overall generalization performance.
+Hybrid design with Local Attention
+RAT is a hierarchical architecture, but not a hybrid model, as it applies the same strategy across all tokens, layers, and heads.
+It is compatible with various hybrid strategies and offers more flexibilities in hybrid modeling by varying the chunk size in different layers or heads.
+In particular, we explore interleaving RAT with sliding-window attention (SWA) [longformer, gau, transformer-xl], a widely adopted technique in recent models [cohere_llm, samba, gateddeltanet, griffin].
+We find that the two are highly complementary: while local attention methods allocate most computation within fixed windows, RAT reserves attention for global access and handles local modeling more efficiently.
+Interleaving them enables the model to efficiently and effectively capture both short-range and long-range dependencies.
 3.2 Efficiency
-We discuss the efficient implementation of RAT for the three main scenarios in language modeling, and provide pseudocode for training or prefilling in 1 and for generation in 2.
-Training or prefilling
-In this setting, a sequence of tokens is processed in parallel, requiring the model to support efficient parallel computation.
-For the intra-chunk recurrent component in Eq.
-(missing) 3, we use PyTorch’s higher order operator associative scan to implement forward and backward passes with $\mathcal{O}(T)$ FLOPs.
-Compared to RNN, RAT applies the recurrent operation to both key and value representations, but chunking can reduce the scan depth from $\mathcal{O}(\log{T})$ to $\mathcal{O}(\log{L})$ and increase parallelism.
-In principle, if the chunk size $L$ is sufficiently small, even non-linear RNNs could be viable in terms of efficiency, which we leave for future work.
-For inter-chunk attention, each token also attends to its own chunk’s key and value, which vary due to causal masking.
-To handle this efficiently, we adopt an online softmax formulation [18]: we separately compute $f({\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top})\tilde{{\bm{V}}}_{:,-1}$ and $f({\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top})\tilde{{\bm{v}}}_{c,l}$, then combine them by adjusting the softmax denominator.
-The first term is computed efficiently and in parallel using PyTorch’s flex attention, which supports custom causal masks and returns the softmax normalization term.
-The second term is implemented via a simple einsum.
-Generation
-In this setting, tokens are generated sequentially.
-The intra-chunk recurrent component can be implemented directly using the recurrence equation, as it only requires a single-step update and does not rely on parallel operators.
-For inter-chunk attention, standard implementations like flash attention [19] can be used without any modification, since no complex causal masking is required in the generation case.
-In summary, the FLOPs per token of RAT is $\mathcal{O}(C\cdot D)$, compared to $\mathcal{O}(D)$ for RNN and $\mathcal{O}(T\cdot D)$ for full attention, where $C$ is the number of chunks, $D$ the model dimension and $T$ the sequence length.
-Notably, our current implementation achieves significant speed-ups in both training and generation without relying on custom CUDA or Triton kernels, though such kernels could further improve performance.
-Parallelism is particularly important for models with more than 7B parameters.
-We believe that RAT is compatible with both tensor and context parallelism.
-In particular, using a chunk size larger than 16 may reduce the communication overhead of ring-style attention in context parallelism.
+We discuss efficiency-related aspects of RAT, and show that our current implementation does not rely on custom CUDA or Triton kernels, yet achieves significant speed-ups in both training and generation in experiments.
+The pseudocode for training is provided in 1, and the decoding algorithm is shown in 2.
+To begin, the FLOPs per token of RAT are $\mathcal{O}(C\cdot D)$, compared to $\mathcal{O}(D)$ for RNN and $\mathcal{O}(T\cdot D)$ for full attention, where $C$ is the number of chunks, $D$ the model dimension, and $T$ the sequence length.
+Most components of RAT are simple and easy to implement, except that the chunk-based design introduces a non-trivial causal masking challenge during training, which we address below.
+Causal masking problem
+In training, where tokens are processed in parallel, special care must be taken to apply causal masking in inter-chunk attention.
+First, a block-wise causal mask is required.
+Second, each token must also attend to its own chunk’s key and value, which should be gated up to its position.
+They vary across tokens due to causal masking, preventing efficient parallel computation.
+To address this, we adopt an online softmax formulation [onlinesoftmax]: we separately compute $f({\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top})\tilde{{\bm{V}}}_{:,-1}$ and $f({\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top})\tilde{{\bm{v}}}_{c,l}$, and then combine the results by adjusting the softmax denominator.
+The first term can be implemented in parallel using existing attention frameworks, while the second is handled via a simple einsum.
+Practical implementation
+For training, we implement intra-chunk recurrence in Eq.
+() 3 using PyTorch’s associative scan, enabling forward and backward passes with $\mathcal{O}(T)$ FLOPs.
+Compared to full-sequence recurrence, chunking reduces scan depth from $\mathcal{O}(\log T)$ to $\mathcal{O}(\log L)$ and thus improves parallelism.
+For inter-chunk attention, we use PyTorch’s flex attention to implement the first term above.
+It supports flash attention with flexible features such as custom masks and returning the softmax denominator, and thus aligns well with our needs.
+For decoding, tokens are generated sequentially.
+Intra-chunk recurrence only requires single-step updates and can be implemented directly.
+For inter-chunk attention, standard implementations like flash attention [flashattention] can be used without modification, as no complex causal masking is required at inference time.
+Parallelism
+We think RAT is compatible with both tensor parallelism and context parallelism, which are commonly used for large model dimensions and long sequence lengths, respectively.
+Both the recurrence and cross-chunk attention in RAT are head-independent, making it easy to apply standard tensor parallelism by assigning different heads to different GPUs.
+For context parallelism, the intra-chunk recurrence is chunk-independent, allowing chunks to be distributed across GPUs.
+Since RAT stores much fewer number of chunk-level key/value vectors (e.g., 16× fewer than full attention), the cross-chunk attention may even avoid ring-style communication.
 4 Experiments
-In this section, we present the efficiency (training, prefilling, and generation) and precision (pretraining, evaluation, and SFT) of different models.
-We focus primarily on the 1.3B model scale with a 4K context window, while additional discussions, including length generalization, 16K context window training, and more evaluation results, are presented in Appendix C.
+In this section, we present the efficiency and large-scale evaluations of RAT, along with comparisons to other models. Additional discussions are provided in Appendix C.
 4.1 Setup
-For brevity, we summarize the setup of the 1.3B model with the 4K context window, which is used in most of our experiments, and provide full details in Appendix A.
-Model.
+For brevity, we summarize the setup for the 1.3B model with a 4K context window, which is used in most of our experiments. Full implementation details are available in Appendix A.
+Model
 We adopt a Transformer architecture that interleaves a token mixing block with a hidden-state mixing block (FFN), each wrapped with residual connections and LayerNorm.
-We compare variants using different token mixing modules, including RNN, attention, and RAT.
-Inspired by recent designs that alternate global and sliding-window attention (SWA), we also explore hybrid models.
-Specifically, Attention-SWA alternates standard attention with SWA (window size 1024) every two layers, while RAT-SWA replaces global attention with RAT.
+We compare variants that use different token mixing modules, including RNN, attention, and RAT, as well as recent state space and linear attention models.
+In addition, we provide hybrid models that interleave RAT with sliding-window attention (window size 1024), such as Attention-SWA and RAT-SWA.
 Efficiency
-We benchmark the latency of a single token mixing block, including input and output projections, on the single H100 in the GH200 system with 120GB of memory.
-For training and prefilling, we report the latency of processing the entire sequence.
-For generation, which is often the most time-consuming case, we report the latency of generating a batch of tokens at a given position.
-To ensure a fair comparison with strong baselines, we use flash attention for the attention module and the associative scan operator for RNN.
-For RAT, we follow our method design as described above.
-All models are wrapped with torch.compile and evaluated in bfloat16 using torch.cuda.amp.
+We benchmark the latency of a single token mixing block, including input and output projections, on a single H100 GPU (GH200 system, 120GB), and also report the maximum throughput of the full model.
+We measure the time required to train on a full sequence or to generate a batch of tokens at specific positions.
+For fair comparison, we use flash attention for the attention baseline and associative scan for RNN.
+All models are compiled using torch.compile and evaluated in bfloat16 with torch.cuda.amp.
 Accuracy
-We pretrain the 1.3B models on a 100B token subset of the FineWeb-Edu dataset [20].
-We use a learning rate of 8.0e-4, decayed to 1.0e-6 using cosine annealing, and a global batch size of 2M tokens.
-Hyperparameters are determined by DeepSeek’s attention-optimized guidelines [21] and are shared across all model variants for fair comparison.
-Then, we further evaluate our models using two types of benchmarks.
-The first is a suite of classical commonsense reasoning tasks [22] which are commonly used to assess the reasoning ability of pre-trained models in short contexts.
-The second is a set of English tasks from LongBench [9], where the average input length for nearly every task exceeds 4,000 tokens, and some tasks have average lengths over 10,000 tokens.
-LongBench tasks often include specific prompts and instructions, which pretrained-only models struggle to follow reliably.
-To better assess the model capability, we also compare different architectures after SFT.
-Specifically, we choose NarrativeQA [23] with two modes, QMSum [24], and WikiSum [25], which vary in sequence length and remain close to the pretraining domain.
-3.2 Efficiency study
-Fig.
-3 presents the efficiency comparison of RNN, RAT, and attention across three scenarios: training, prefilling, and generation.
-For training and prefilling, when the sequence length is short (e.g., 4096), RAT is slightly slower than attention due to under-utilized GPU parallelism in the flex attention component (with only 256 chunks) and the overhead of associative scan.
-We believe this can be improved with better CUDA kernel optimization.
-As the sequence length increases, RAT becomes more efficient: for $L=16$, we observe about 2×2$\times$2 × speedup at 16K, 3×3$\times$3 × at 32K, 4×4$\times$4 × at 64K, and 7×7$\times$7 × at 100K tokens.
-Additional results in Subsec.
-B.1 show up to 910×9-10$\times$9 - 10 × at 200K.
-While RNN slows down beyond 32K, RAT benefits from increased intra-chunk parallelism and retains an advantage over attention.
-For generations, RAT(L=16) achieves 9×9$\times$9 × speedup at position 4K and around 10×10$\times$10 × for longer sequences.
-It also reduces the usage of the KV cache, making it much less prone to OOM.
-RAT(L=64) can be up to 4×4$\times$4 × faster than $L=16$ with larger batch sizes or longer inputs.
-Figure 3: Latency results.
-Latency of a single RNN, RAT, or attention layer (including input and output projections) with model dimension of 2048 under (a) training, (b) prefilling, and (c) generation.
-(a,b) report full-sequence latency on 200K tokens; (c) reports latency for generating 512 tokens at a fixed position.
-RAT also reduces KV cache usage, lowering OOM risk.
-See Appendix B for more results.
-3.3 Ablation study
+We pretrain all 1.3B models on 100B tokens from FineWeb-Edu [fineweb_edu] using the same setup: a learning rate of 8.0e-4 decayed to 1.0e-6 (cosine schedule) and a global batch size of 2M tokens, following DeepSeek’s hyperparameter guidelines [deepseek_hp].
+We directly evaluate the models on both short-context tasks—several classical commonsense reasoning benchmarks [eval-harness]—and long-context tasks, including 11 tasks from LongBench [longbench] covering QA, summarization, and code completion.
+Since LongBench includes instruction-heavy prompts that pretrained models often struggle with, we also include SFT-based evaluations.
+Specifically, we use NarrativeQA [narrativeqa] (two modes), QMSum [qmsum], and WikiSum [wikisum] to test long-context understanding with SFT.
+To assess retrieval capabilities, we include the Ruler benchmark [ruler] and test nine synthetic needle-in-haystack tasks with varying configurations.
+A single round of lightweight fine-tuning is applied to adapt models to specific prompts.
+4.2 Analyses of RAT
+Figure 2: Latency of the temporal mixing block (including linear projections) with a model dimension of 2048.
+(a): full-sequence latency with 200K tokens; (b): generation of 512 tokens at specified positions.
+We adopt flash attention for Attention.
+Table 2: Maximum throughput of full models (tokens/sec), measured by generating 1024 tokens from a 3072-token prompt.
+By reducing the KV cache memory and boosting speed, we achieve 10× maximum throughput compared to flash attention , and even more on 13B models, as attention suffers from poor GPU utilization at larger scale.
+[…]
+Efficiency study
+Subsec.
+4.2 and Table 2 present the efficiency comparison, with additional results provided in Appendix B.
+For training, on the strong H100 GPU, when the sequence length is short (e.g., 4096), RAT is slightly slower than attention due to underutilized GPU parallelism in the flex attention (with only 256 chunks) and the overhead introduced by associative scan.
+We expect these can be improved through further kernel-level optimizations.
+As the sequence length increases, RAT becomes increasingly efficient: for $L=16$, we observe approximately $2\times$ speedup at 16K, $3\times$ at 32K, $4\times$ at 64K, and $7\times$ at 100K tokens.
+For generation, RAT(L=16) achieves a $9\times$ speedup at position 4K and around $10\times$ for longer sequences.
+It also reduces KV cache usage, making it significantly less prone to out-of-memory (OOM) errors and enabling much higher maximum throughput.
+Ablation study
 We conduct the ablation study on a 200M-parameter model trained on the book dataset.
@@
 Figure 3: (a) Ablation study on RAT(L=64). (b) and (c) show pretraining results on 200M and 1.3B models, respectively. RAT lies between RNN and attention in terms of pretraining perplexity.
-3.4 Pretraining
-We begin our study by examining pretraining perplexity as a foundational measure of architectural performance, showing that RAT lies between attention and RNN.
-The main results are presented in Fig.
-4(b) and (c).
-In (b), on a 200M model, increasing the size of the chunk leads to higher perplexity.
+Pretraining
+In Fig.
+3(b) and (c), we begin our study by examining pretraining perplexity as a foundational measure of architectural performance, demonstrating that RAT falls between attention and RNN.
+In (b), using a 200M model, we observe that increasing the chunk size leads to higher perplexity while reducing FLOPs.
 Viewing attention as $L=1$ and RNN as $L=T$, we observe perplexities of 13.26 for attention, 13.42 for RAT(L=64), 13.72 for RAT(L=256), and 14.31 for RNN.
-Interestingly, increasing the training context from $T=8192$ to $T=32768$ causes a sharp increase in perplexity for RNN, while attention and RAT remain stable.
-Note that on a 200M model, the PPL may not decrease steadily when going to longer sequences.
-We then scale up to a 1.3B model trained on web data, which primarily consists of short sequences (often under 2K tokens) concatenated with separator tokens to construct 4K-length contexts.
-In this regime, we find that the chunk length should be reduced to 16 to match the performance of attention.
-Once again, increasing the size of the chunks and thus reducing the FLOPs leads to higher perplexity, strengthening the trade-off and intermediate nature of RAT.
-3.5 Downstream evaluation
-Table 2: We evaluate our model on six commonsense reasoning tasks using the lm-evaluation-harness [22].
-These tasks have very short contexts, with each sample containing no more than 300 tokens.
-Normalized accuracy is reported when available, except for WinoGrande and LAMBADA, which use standard accuracy.
-Model ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg.
-RNN 38.57 66.46 55.95 39.74 72.69 53.99 54.57
-RAT(L=16) 39.33 67.80 56.44 44.40 72.69 53.20 55.64
-Attention 37.71 66.67 57.44 47.84 72.91 57.70 56.71
-Attention-SWA 37.20 66.20 57.09 47.97 72.74 56.91 56.35
-RAT(L=16)-SWA 40.78 68.22 57.85 49.33 73.94 56.91 57.84
+Interestingly, increasing the training context from $T=8192$ to $T=32768$ leads to a sharp rise in perplexity for RNN, while both attention and RAT remain stable, demonstrating the pitfalls of full-sequence compression.
+We then scale up to a 1.3B model trained on web data, which primarily consists of short sequences (often under 1K tokens) concatenated with separator tokens to construct 4K-length contexts.
+We find that RAT(L=16) gives the comparable performance to attention, and again RAT with different chunk sizes exhibits behavior intermediate between attention and RNN.
+4.3 Large-scale evaluation and comparisons
+In this section, we present direct evaluation and SFT results of our trained 1.3B models.
+We mainly compare RAT(L=16) with its two extremes: RNN and attention.
+Additionally, we include results from recent state space and linear attention models [gateddeltanet], which are trained under the same settings (model size, sequence length, and dataset) as ours.
+We report the value/slot sizes, which indicates the size of memory slots and FLOPs, based on sequence length $T$, model dimension $D$, and number of layers $N$.
+Note that state space and linear attention models use fixed-size memory and expand the classic recurrence state, while RAT scales its memory capacity with the sequence length.
 Commonsense reasoning
-The performance gap between RNN and attention is small on short and relatively simple commonsense reasoning tasks, typically within two points on average.
-This suggests that RNN can be surprisingly competitive in short-context settings.
-RAT(L=16) consistently outperforms RNN in most tasks, with particularly strong gains on LAMBADA, highlighting its advantage in modeling token dependencies.
-However, it performs poorly on WinoGrande compared to attention.
-One possible explanation is that WinoGrande relies heavily on subtle sentence-level cues and fine-grained referential resolution; moreover, the variance may also stem from the instability of the task itself.
-Interestingly, interleaving RAT with SWA yields a hybrid model (RAT(L=16)-SWA) that achieves the best overall performance in almost all tasks.
+In Table 3, the performance gap (within two points on average) between RNN and attention is small on these short commonsense reasoning tasks, indicating that RNN can be surprisingly competitive in short-context settings.
+RAT(L=16) consistently outperforms RNN on most tasks and surpasses other recent models on 4 out of 7 benchmarks, with only at most $20ND$ memory slots.
+We also observe that some binary-choice tasks (e.g., BoolQ, WinoGrande) show inconsistent results under small-scale pretraining.
+For instance, attention struggles on BoolQ, while RAT underperforms on WinoGrande.
+Interestingly, the hybrid variant RAT(L=16)-SWA, which interleaves RAT with SWA, achieves the best overall accuracy across almost all tasks.
 This suggests a complementary architecture: using SWA for strong local attention and RAT for long-range dependencies can bring us a model that is not only more efficient but also more accurate than its dense counterpart.
-In addition, the hybrid version mitigates the weaknesses of RAT in WinoGrande, improving overall robustness.
-We do not include results from Gated DeltaNet [26], a recent state space model that also interleaves with sliding-window attention, [^1] While our best model outperforms their best model (e.g., 49.33 vs.
-48.76 on LAMBADA, 73.94 vs.
-72.57 on PIQA, and 57.85 vs.
-56.88 on HellaSwag), we do not include their results in our tables.
-Their approach follows a fundamentally different modeling paradigm—state space modeling—whereas our work focuses on an intermediate design between recurrence and attention.
-These differing motivations, together with discrepancies in window size (their 2048 vs.
-our 1024) and training protocols, make direct comparisons less meaningful.
-as their modeling paradigm and experimental setup differ substantially from ours.
-Table 3: Evaluation results on LongBench.
-For each task, we highlight the top two values to illustrate general trends on instruction-based tasks.
-To save space, task numbering follows the original LongBench paper; the number-to-task mapping and token counts (measured with the LLaMA2 tokenizer) are provided in Subsec.
-C.4.
+Table 3: We evaluate our models with model dimension $D=2048$ and number of layers $N=24$ on seven commonsense reasoning tasks using lm-evaluation-harness [eval-harness].
+Since all tasks (except BoolQ, which has only 13 sequences >300>300 tokens) have sequences $\leq 300$ tokens, we set $T=300$.
+Results marked with * are from a recent study [gateddeltanet] trained on the same setting as ours.
+For SWA, slot sizes remain $150ND$ due to the short context.
+Model Value/Slot size ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande BoolQ
+$T=300$ acc/acc_n acc/acc_n acc/acc_n acc acc/acc_n acc acc
+RNN NDND 35.15/38.57 70.66/66.46 42.43/55.95 39.74 71.49/72.69 53.99 62.54
+Mamba* $64ND$ -/35.40 69.52/- -/52.91 -/43.98 71.32/- 52.95 61.13
+Mamba2* $256ND$ -/37.88 72.47/- -/55.67 -/45.66 71.87/- 55.24 60.13
+DeltaNet* $128ND$ -/35.66 68.47/- -/50.93 -/42.46 70.72/- 53.35 55.29
+GatedDeltaNet* $288ND$ -/38.39 71.21/- -/55.76 -/46.65 72.25/- 57.45 60.24
+Attention $300ND$ 35.67/37.71 71.25/66.67 44.16/57.44 47.84 73.29/72.91 57.70 58.23
+RAT(L=16) $\sim 20ND$ 36.35/39.33 72.64/67.80 43.27/56.44 44.40 72.03/72.69 53.20 63.30
+GatedDeltaNet-SWA* 288N/2D300N/2D288N/2D-300N/2D -/40.10 71.75/- -/56.53 47.73 72.57/- 58.4 63.21
+Attention-SWA 300N/2D300N/2D300N/2D-300N/2D 35.24/37.20 71.97/66.20 44.36/57.09 47.97 72.25/72.74 56.91 61.25
+RAT(L=16)-SWA $\sim 20N/2D-300N/2D$ 37.20/40.78 72.56/68.22 44.33/57.85 49.33 73.29/73.94 56.91 63.21
+Table 4: Evaluation results on LongBench with $T=4096$, $D=2048$, and $N=24$.
+NQA: NarrativeQA, MQA: MultiFieldQA-en, HQA: HotpotQA, WQA: 2WikiMultihopQA, MSQ: Musique, GR: GovReport, MN: MultiNews, TQA: TriviaQA, RBP: RepoBench-P.
 Metrics: F1 for QA tasks, Rouge-L for summarization, and EditSum for code tasks.
-Specifically, accuracy is used for 4-1, F1 for 4-2, and Rouge-L for 4-3.
-Model Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
-1-1 1-2 1-3 Avg. 2-1 2-2 2-3 Avg. 3-1 3-2 3-3 Avg. 4-1 4-2 4-3 Avg. 6-1 6-2 Avg.
-RNN 11.7 8.8 19.5 13.3 9.1 15.1 5.4 9.9 16.9 16.8 16.4 16.7 10.0 17.5 13.6 13.7 14.0 17.6 15.8
-RAT(L=64) 13.9 13.0 27.0 18.0 14.9 16.3 4.3 11.8 18.9 17.7 19.9 18.8 17.5 19.2 15.2 17.3 16.2 19.7 18.0
-RAT(L=16) 14.5 16.1 28.3 19.6 16.7 18.9 6.3 14.0 24.9 17.5 18.3 20.2 38.5 18.2 18.7 25.1 14.2 20.6 17.4
-Attention 12.3 14.0 28.2 18.2 13.4 17.7 4.9 12.0 18.6 18.9 21.0 19.5 39.5 19.6 23.9 27.7 21.3 26.5 23.9
-Attention-SWA 13.1 14.6 24.5 17.4 13.7 19.0 6.1 12.9 19.4 17.3 21.4 19.4 47.0 19.6 25.1 30.6 18.7 24.6 21.7
-RAT(L=16)-SWA 12.7 15.6 28.2 18.8 14.2 18.1 7.4 13.2 20.1 18.6 19.8 19.5 33.5 18.4 25.6 25.8 26.3 30.1 28.2
+Results marked with * are from a recent study [gateddeltanet] trained under the same settings as ours.
+Note that they use local attention with a window size of 2048, while we use 1024.
+Model Value/Slot size Single-Document QA Multi-Document QA Summarization Code Completion
+$T=4096$ NQA Qasper MQA Avg. HQA WQA MSQ Avg. GR QMSum MN Avg. LCC RBP Avg.
+RNN NDND 11.7 8.8 19.5 13.3 9.1 15.1 5.4 9.9 16.9 16.8 16.4 16.7 14.0 17.6 15.8
+Mamba2* $256ND$ 11.1 11.3 18.6 13.7 11.8 15.1 6.7 11.2 6.7 14.5 7.4 9.5 17.9 20.6 19.3
+DeltaNet* $128ND$ 12.9 10.8 21.5 15.1 10.9 13.2 5.1 9.7 6.5 13.5 7.2 9.1 17.6 20.3 19.0
+GatedDeltaNet* $288ND$ 14.1 14.0 23.3 17.1 13.7 14.4 5.8 11.3 7.5 16.4 7.9 10.6 18.7 22.1 20.4
+Attention $4096ND$ 12.3 14.0 28.2 18.2 13.4 17.7 4.9 12.0 18.6 18.9 21.0 19.5 21.3 26.5 23.9
+RAT(L=64) $64ND$ 13.9 13.0 27.0 18.0 14.9 16.3 4.3 11.8 18.9 17.7 19.9 18.8 16.2 19.7 18.0
+RAT(L=16) $256ND$ 14.5 16.1 28.3 19.6 16.7 18.9 6.3 14.0 24.9 17.5 18.3 20.2 14.2 20.6 17.4
+GatedDeltaNet-SWA* 288N/2D2048N/2D288N/2D-2048N/2D 14.5 12.3 26.6 17.8 12.6 23.6 6.1 14.1 9.1 16.1 12.8 12.7 15.5 19.2 17.4
+Attention-SWA 4096N/2D1024N/2D4096N/2D-1024N/2D 13.1 14.6 24.5 17.4 13.7 19.0 6.1 12.9 19.4 17.3 21.4 19.4 18.7 24.6 21.7
+RAT(L=16)-SWA 256N/2D1024N/2D256N/2D-1024N/2D 12.7 15.6 28.2 18.8 14.2 18.1 7.4 13.2 20.1 18.6 19.8 19.5 26.3 30.1 28.2
 LongBench
-We present the results of the evaluation for 14 English tasks in LongBench in LABEL:tab:longbench_4k.
-It is expected that no single model consistently dominates across all tasks, since LongBench employs heavy instruction-style prompts and the tasks span diverse domains.
-Evaluating pretrained-only models without any fine-tuning on these instruction-based tasks remains challenging; even in the original paper, 7B pretrained-only models perform much worse than tuned ones.
-Therefore, we focus on general performance trends across tasks, particularly average scores by task category and the top-2 models for each task.
-First, RAT(L=16) and its hybrid variant RAT(L=16)-SWA consistently rank among the top two models in nearly all task groups.
-This is notable given that RAT(L=16) lags behind attention in pretraining perplexity and performs slightly worse in short-context commonsense reasoning.
+In LABEL:tab:longbench_4k, we present results on 11 long-context tasks from LongBench.
+Since LongBench includes instruction-heavy prompts across diverse domains, it’s natural that no single model dominates all tasks.
+Evaluating purely pretrained models on such instruction-based datasets can be very challenging.
+Therefore, we focus on general performance trends, especially average scores by task category.
+As shown in the table, RAT(L=16) and its hybrid variant RAT(L=16)-SWA achieve top performance in many groups, such as question answering and summarization.
+This is noteworthy, given that RAT(L=16) slightly lags behind attention in pretraining perplexity and underperforms in short-context commonsense reasoning.
 We hypothesize that its advantage on LongBench stems from structural adaptation to very long sequences—for instance, achieving 24.9 on GovReport (3-1) (vs.
@@
 13.4).
-Second, unlike the small gap ( 2 points) between RNN and attention on short-context tasks, the difference becomes much larger on LongBench, reaching up to 8 points on code completion.
-Finally, we note some outliers: RAT-based models underperform on TREC (4-1), a few-shot QA task evaluated via classification accuracy, for reasons not yet fully understood.
-In addition, RAT(L=16) performs poorly on code completion.
-However, its hybrid variant RAT(L=16)-SWA significantly improves in this setting, outperforming the second-best model by more than four points, consistent with earlier observations on commonsense reasoning tasks.
-3.6 Supervised Fine-tuning
-Since LongBench tasks are instruction-based and pretrained-only models often struggle to follow specific prompts, we further assess the performance of SFT on the QA and summarization tasks in the long-context setting, as shown in Table 4.
-For NarrativeQA, where passages are extremely long and may not contain the answer within the truncated 4K-token context, we evaluate two settings: (1) summary-only, which uses only the passage summary; and (2) summary+passage, which concatenates the summary and full passage.
-While the latter provides more context, it also introduces much noise.
+Second, unlike the small gap ($\sim$2 points) between RNN and attention on short-context tasks, the difference becomes much larger on LongBench, further demonstrating the problem of full-sequence compression.
+We also notice that RAT(L=16) performs less favorably on code completion.
+However, its hybrid variant significantly improves performance in this setting, outperforming the second-best model by about four points.
+Supervised Fine-tuning
+Since LongBench tasks are instruction-based and pretrained-only models often struggle to follow specific prompts, we further assess the performance of SFT on some long-context QA and summarization tasks, as shown in Table 5.
+For NarrativeQA, where passages are extremely long and may not contain the answer within the truncated 4K-token context, we evaluate two settings: (1) summary-only, which uses only the passage summary, and (2) summary plus passage, which concatenates the summary with the full passage.
 In both cases, RAT(L=16)-SWA delivers strong performance while remaining efficient.
-Although the summary and passage contain useful information, we observe that models, particularly the attention module, perform worse in the second setting, with F1 dropping to 33.
-This highlights the difficulty in identifying and focusing on relevant content in noisy, extended contexts.
-In this setup, attention can be more easily distracted due to the absence of structural bias, which can lead to flattened attention score distributions [27].
-In contrast, models based on RAT appear to be more robust, likely due to their chunk-wise structure.
+Although both the summary and passage contain useful information, we observe that models, particularly the attention-based ones, perform worse in the second setting, with F1 dropping to 33.
+We hypothesize that in this setup, attention may be more easily distracted due to the absence of structural bias, which can lead to flattened attention score distributions [scalablesoftmax].
+In contrast, models based on RAT appear more robust, likely due to their chunk-wise structure.
 For summarization tasks, RNN performs better than in QA, and RAT(L=16)-SWA achieves the best result on QMSum, surpassing the next-best model by approximately 1 Rouge-L point.
@@
 […]
-[…]
+Model Value/Slot size $^{\textbf{1}} NarrativeQA$$^{\textbf{2}}$ QMSum WikiSum RNN $ND$ 30.5 27.4 23.2 34.3 RAT(L=16) $256ND$ 60.8 43.5 23.3 36.6 Attention $4096ND$ 61.3 33.0 23.4 36.8 Attention-SWA 4096N/2D1024N/2D4096N/2D-1024N/2D 63.3 39.6 23.4 36.7 RAT(L=16)-SWA 256N/2D1024N/2D256N/2D-1024N/2D 63.2 43.7 24.6 36.7
+Table 6: Retrieval ability: Accuracy performance with exact match scoring on the Needle-in-Haystack tasks with different configurations from the RULER benchmark [ruler].
+We use $T=4096$, $D=2048$, and $N=24$.
+Model Value/Slot size single_1 single_2 single_3 multikey_1 multikey_2 multikey_3 multiquery multivalue
+RNN NDND 99.4 99.8 0.0 2.2 0.0 0.0 5.8 5.7
+RAT(L=64) $64ND$ 100.0 99.8 55.6 96.4 62.2 1.0 86.3 89.1
+RAT(L=16) $256ND$ 99.6 100.0 94.6 99.6 99.6 82.6 91.2 94.8
+Attention $4096ND$ 100.0 100.0 100.0 99.6 99.6 95.2 98.6 99.1
+Retrieval ability
+In Table 6, we evaluate the three models on challenging synthetic needle-in-haystack tasks that test retrieval ability (see Subsec.
+C.3 for task descriptions).
+It can be seen that attention consistently performs well due to its full-token access.
+RNN handles simpler tasks like single_1 and single_2 reasonably well, but its performance drops on harder settings.
+RAT(L=16) matches attention closely on most tasks but struggles with UUID ones (single_3, multikey_3) due to their difficulty and evaluation metric as exact match scoring.
+Meanwhile, RAT(L=64) falls between RNNs and RAT(L=16), as expected given its partial access to long-range context.
+These results align with the model structures: attention sees all tokens directly, recurrence compresses all history into a single state, and RAT blends both, preserving chunk-level access while reducing FLOPs.
 5 Related work
-RNN
-Recurrent neural networks (RNNs) have a long history of modeling sequential data [28, 29, 30], but suffer from training instabilities and lack parallelizability, especially when handling long sequences.
-Various approaches have been proposed to improve their trainability, such as gating mechanisms [10, 11], unitary designs [31, 32], and partially parallelizable variants [33, 34].
-Moreover, RNNs face a memory bottleneck as they compress entire sequences into fixed-size hidden-state vectors.
-Some earlier work addressed this issue by allowing the current step to access earlier time steps directly [35, 36, 37].
-For example, Koutnik et al.
-[36] introduces multiple timescales by updating hidden states only at predefined clock rates, while Ke et al.
-[37] proposes sparse, learnable connections to past tokens, with weights computed by a small MLP.
-More recently, state space models [38, 12, 39, 13, 40] have emerged, expanding hidden states from vectors to matrices through state expansion and reduction techniques.
-They also replace standard matrix multiplications with structured operations (e.g.
-diagonal matrices) to further enable parallel computation [14].
-In contrast to these approaches, and inspired by the success of attention mechanisms, we propose integrating attention directly over the RNN history to allow explicit retrieval of past information.
-This design remains simple and naturally interpolates between RNN and attention behaviors.
-By applying inter-chunk attention to retrieve information, we avoid state expansion, which may introduce training instability and typically requires careful normalization or initialization [12, 5].
-Attention
-While attention mechanisms have proven highly effective, they suffer from slow computation due to their full access to all previous tokens, especially in long sequences.
-Recently, there has been a growing interest in linear attention [4, 6, 41, 13, 5, 42, 26], which explores the dual formulation of linear RNN and linear attention.
-These approaches rely on strict linearity constraints and introduce state expansion and per-head gating techniques.
-In contrast to this duality perspective, we adopt an intermediate view between per-dimension-gated RNNs and softmax-based attention, without imposing hard linearity constraints on either side.
-The use of linear recurrence in our model is primarily motivated by training efficiency, but it is not strictly required at the design level.
-Other works have explored attention within local contexts, often motivated by the diagonal structures observed in attention maps, and have introduced recurrence or sparse attention mechanisms to provide access to earlier context [1, 2, 43, 44, 3, 45].
-For example, Ma et al.
-[43] applies attention in chunks after gating the query and key vectors, and Behrouz et al.
-[3], inspired by cognitive science, focuses attention on the short-range context.
-In our work, we take a different approach: we use attention to model long-range dependencies, leveraging their strength in directly accessing distant tokens, while applying RNNs to short-range contexts, where they are typically less prone to memory degradation.
-This can also be seen as a form of attention with resolution, loosely mirroring how humans process information: short-term inputs are integrated into coherent representations, while long-term events are stored as memory anchors that can be selectively retrieved when relevant.
-We also find that RAT and the attention of sliding windows [1] are structurally complementary.
-The interleaving of the two leads to further performance improvements while being efficient.
+State space and linear attention models
+Recurrent neural networks (RNNs) have long been used for sequential modeling [elman1990finding, siegelmann1992computational, bengio1994learning].
+Recent state space models [s4, s5, mamba, griffin, lru] extend recurrence by expanding hidden states from vectors to matrices for higher accuracy, removing non-linearity and using structured operations (e.g., diagonal matrices) to enable parallel computation [s5].
+Per-head gating connects these models [mamba2] to linear attention [linear_attention, retnet, rwkv, lru, deltanet, gateddeltanet] and shows the duality between linear recurrence and linear attention.
+Unlike this duality perspective, we take an intermediate view between per-dimension-gated RNNs and softmax-based attention, without enforcing strict linearity.
+Linear recurrence in RAT is intended primarily for training efficiency, rather than imposed as a design constraint.
+Despite state expansion, these models inherit the fixed-size representation limitation of classic RNNs, leading to degraded memory on long sequences.
+A concurrent work [log_linear_attention] proposes addressing this by growing the memory slots logarithmically.
+We similarly target the fixed-size memory problem, but use softmax-based attention to access distant tokens.
+Hybrid strategies [samba, hymda] combining state space models and attention are also popular.
+While RAT is not a hybrid with the same computation graph for all tokens and all layers, it is orthogonal to such designs and provides more flexibility, such as using different chunk sizes for different layers and heads, which we leave for future work.
+Softmax-based attention
+Attention mechanisms suffer from slow computation due to their full-token access.
+Earlier works have explored local attention with chunking, combined with recurrence or sparse mechanisms to access earlier context [longformer, gau, mega, megalodon, landmarkattention].
+For example, mega gates query/key vectors before local attention, while gau adds the outputs of local attention to cross-chunk linear attention.
+In our work, we instead leverage attention’s strength in directly accessing distant tokens with lightweight recurrence for local contexts, and hierarchically organize their outputs.
+As shown in Subsec.
+C.1, RAT utilizes FLOPs more effectively than local attention.
+We also regard RAT and local attention as structurally complementary.
+Although we experimented with interleaving them in this work, we leave the combination with more advanced local attention variants in future research.
+RAT can be viewed as an attention mechanism with resolution, akin to how humans process information: short-term inputs are integrated into coherent representations, while long-term events are stored as memory anchors for selective retrieval.
+Compared to dilated attention [longformer], our simple recurrence enables global perception, and the chunk-level attention has high training parallelism and reduced cache memory (see Subsec.
+C.1 for more).
+Finally, a recent work [gatedslotattention] also aggregates the key and value vectors and applies softmax-based attention over them.
+However, it compresses the full-sequence vectors into a 2D memory matrix, unlike our chunk-based design, which enables flexible memory slots.
 6 Conclusion
@@
 But as shown in Subsec.
-C.2, this can be significantly reduce by shorter inter-chunk attention spans, and we also explore the use of NoPE.
+C.2, this can be significantly reduced by shorter inter-chunk attention spans, and we also explore the use of NoPE.
+Acknowledgments and Disclosure of Funding
+We sincerely thank the RCP and the SCITAS team at EPFL for GPU support.
+We also thank the Swiss AI Initiative and the Swiss National Supercomputing Centre (CSCS) for supporting this work through grants under project IDs a06 and a10.
+We thank Skander Moalla, Anja Surina, and Lars Quaedvlieg for helpful discussion on SFT tasks.
+We extend our appreciation to Karin Getaz for administrative support.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We ensure that the abstract and introduction clearly summarize the paper’s core contributions: the RAT architecture, its intra- and inter-chunk mechanisms, the efficiency, and empirical evaluations on both pretraining and downstream tasks.
+All major claims are supported by experiments and analyses throughout the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss limitations in Sec. 6
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We describe our experimental setup and implementation details in Sec. 4 and Appendix A.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We open-source our code with link put in Appendix. The dataset we use is publicly available.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We describe our experimental setup in Sec. 4 and give details in Appendix A.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34.
+Justification: We do not report error bars or statistical significance due to computational cost and limited variance observed across repeated runs in preliminary testing.
+We follow the common practice in similar-scale LLM experiments.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Computing resources are discussed in Sec. 4 and Appendix A.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Societal impacts are discussed in Appendix D.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54.
+Justification: We do not include safeguards in the current version, as model and data release is not the focus of this work, and the model is only 1B scale.
+Should we release models in the future, we will follow community standards to mitigate misuse risk.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All the assets used have been cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provide documented code.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: LLMs were not used in any component of the core research methodology. Any usage was limited to writing assistance and did not affect the technical contributions of the work.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Implementation details
@@
 Dataset
-In our preliminary studies, we use the PG19 dataset [46], a long-form English book corpus with inherently long contexts.
-For the 1.3B model experiments, we adopt the FineWeb-Edu dataset [20], using its 100B-token randomly sampled version downloaded from the HuggingFace repository.
+In our preliminary studies, we use the PG19 dataset [pg19], a long-form English book corpus with inherently long contexts.
+For the 1.3B model experiments, we adopt the FineWeb-Edu dataset [fineweb_edu], using its 100B-token randomly sampled version downloaded from the HuggingFace repository.
 To match the pretraining context length, we concatenate documents using a separator token.
 Note that web samples are usually very short, compared to the book dataset.
-For downstream evaluation, we consider a suite of classical commonsense reasoning benchmarks from the Eleuther AI evaluation harness [22], including PIQA [47], ARC-C [48], and HellaSwag [49].
+For downstream evaluation, we consider a suite of classical commonsense reasoning benchmarks from the Eleuther AI evaluation harness [eval-harness], including PIQA [piqa], ARC-C [arc], and HellaSwag [hellaswag].
 In the LLaMA2 tokenizer, we observe that the inputs of these tasks typically contain fewer than 300 tokens.
-For the LongBench evaluation, in LABEL:tab:longbench we provide a mapping of task numbering to names and annotate their input lengths for each task, offering a rough indication of task difficulty with respect to input length.
+For the LongBench evaluation, in LABEL:tab:longbench we provide input length for each task, offering a rough indication of task difficulty with respect to input length.
 For SFT-based tasks, we have elaborated sequence lengths in the main text.
 As LongBench and SFT tasks often involve very long inputs, we apply truncation in the middle to preserve information at both the beginning and the end.
-Table 5: We provide a mapping from task numbers to task names.
+Table 7: We report the average input length and the 95th percentile input length of each task, measured in tokens using the LLaMA2 tokenizer.
 NQA (NarrativeQA), MQA (MultiFieldQA-en), HQA (HotpotQA), WQA (2WikiMultihopQA), MSQ (Musique), GR (GovReport), MN (MultiNews), TQA (TriviaQA), and RBP (RepoBench-P).
-We also report the average input length and the 95th percentile input length of each task, measured in tokens using the LLaMA2 tokenizer.
-Task Single-Document QA Multi-Document QA Summarization Few-shot learning Code Completion
-1-1 1-2 1-3 2-1 2-2 2-3 3-1 3-2 3-3 4-1 4-2 4-3 6-1 6-2
-Name NQA Qasper MQA HQA WQA MSQ GR QMSum MN TREC TQA SAMSum LCC RBP
-Avg. 36037 5780 8115 15329 8483 18555 12280 15980 3156 7792 14098 11168 4307 14818
-95th pctl. 77966 10164 14994 19755 16939 20066 25721 29069 7135 12532 25006 19823 10401 31937
+Task Single-Document QA Multi-Document QA Summarization Code Completion
+Name NQA Qasper MQA HQA WQA MSQ GR QMSum MN LCC RBP
+Avg. 36037 5780 8115 15329 8483 18555 12280 15980 3156 4307 14818
+95th pctl. 77966 10164 14994 19755 16939 20066 25721 29069 7135 10401 31937
 200M model
@@
 We use the GPT2 tokenizer.
-Following Mohtashami and Jaggi [45], we repeat the PG19 training split five times to reach a total of 15B training tokens.
+Following landmarkattention, we repeat the PG19 training split five times to reach a total of 15B training tokens.
 The learning rate is scheduled using cosine annealing, starting at $6.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with a warm-up ratio of 10%.
@@
 1.3B model
-The 1.3B-parameter model uses a model dimension of 2048, 24 transformer layers, and a head dimension of 128, equipped with RMSNorm [50] and without bias.
+The 1.3B-parameter model uses a model dimension of 2048, 24 transformer layers, and a head dimension of 128, equipped with RMSNorm [rmsnorm] and without bias.
 The RoPE base is also set to 10,000.
@@
 For pretraining, we use a cosine-annealed learning rate schedule starting at $8.0\times 10^{-4}$ and decaying to $1.0\times 10^{-6}$, with 5% warmup.
-The global batch size is set to 2M tokens.
-We primarily train with a context window of 4K tokens and include additional experiments with a 16K context window, for which the RoPE base is increased to 500,000.
+The global batch size is set to 2M tokens, and the context window is set to 4096.
 Each model is trained on 16 H100 GPUs, requiring approximately 2 to 3 days to complete.
@@
 Although we explored different hyperparameters during the early experimentation, we observed that the relative trends remained largely stable.
-Thus, we fix the learning rate and batch size to $(1.0\times 10^{-5},128)$ for large datasets, and $(1.0\times 10^{-5},32)$ for the smaller QMSum [24] task, following common practice for 1B-scale models.
+Thus, we fix the learning rate and batch size to $(1.0\times 10^{-5},128)$ for large datasets, and $(1.0\times 10^{-5},32)$ for the smaller QMSum [qmsum] task, following common practice for 1B-scale models.
 The weight decay is set as 0.01, and all other hyperparameters follow the pretraining setup.
@@
 B.1 Latency
-To supplement Fig. 3 in main text, we put the concrete latency number of a single layer in Table 6, Table 7, and Table 8.
-Table 6: Single layer (including input and output projections) training time across different sequence lengths. The latency (ms) is tested on 200K tokens.
-Model 4096 8192 16384 32768 65536 131072 262144 RNN 77.00 78.42 77.50 83.50 115.83 130.62 195.30 RAT(L=128) 150.19 150.76 156.65 154.52 159.94 165.07 206.46 RAT(L=64) 146.25 148.46 151.10 155.83 165.58 177.78 227.76 RAT(L=16) 150.08 153.63 158.98 173.89 202.00 249.79 378.11 Attention 122.06 176.44 277.61 474.90 877.14 1713.82 3417.48 Attention/RAT(L=16) 0.81$\times$ 1.15$\times$ 1.75$\times$ 2.73$\times$ 4.34$\times$ 6.86$\times$ 9.04$\times$
-Table 7: Single layer (including input and output projections) prefilling time across different sequences lengths. The latency (ms) is tested on 200K tokens.
-Model 4096 8192 16384 32768 65536 131072 262144 RNN 24.93 24.75 25.32 27.09 30.54 39.40 56.16 RAT(L=128) 44.74 44.78 45.09 45.50 46.67 48.62 51.75 RAT(L=64) 43.46 44.27 44.27 44.83 47.08 50.83 57.38 RAT(L=16) 44.03 44.90 45.53 51.08 57.31 71.18 99.53 Attention 36.93 52.71 80.21 135.68 245.14 494.62 997.50 Attention/RAT(L=16) 0.84$\times$ 1.17$\times$ 1.76$\times$ 2.66$\times$ 4.28$\times$ 6.95$\times$ 10.02$\times$
-Table 8: Single layer (including input and output projections) generation time at the specified position. The latency (ms) is tested on generating batches of tokens with $B=64$, $B=512$, and $B=1024$.
+To supplement Subsec. 4.2 in main text, we put the concrete latency number of a single layer in Table 8, Table 9, and Table 10.
+Table 8: Single temporal-mixing layer (including input and output projections) training time across different sequence lengths. The latency (ms) is tested on 200K tokens.
+[…]
+Table 9: Single temporal-mixing layer (including input and output projections) prefilling time across different sequences lengths. The latency (ms) is tested on 200K tokens.
+[…]
+Figure 4: We measure the maximum throughput of the full 1.3B model for generating 1024 tokens under different prefilling lengths.
+For each total sequence length $T$, the prefilling length is set to T1024T-1024.
+For example, $T=4096$ corresponds to a prefilling of 3072 tokens, while $T=8192$ and $T=16384$ correspond to 7168 and 15360 tokens, respectively.
+When the sequence length increases, the maximum throughput ratio between RAT and attention rises from $10.2\times$ to $15.6\times$, highlighting the strong efficiency advantage of RAT in long-context generation.
+Table 10: Single temporal-mixing layer (including input and output projections) generation time at the specified position.
+The latency (ms) is tested on generating batches of tokens with $B=64$, $B=512$, and $B=1024$.
 Model 4096 8192 16384 32768 65536 131072 262144 $B=64$ RNN 0.36 0.33 0.34 0.33 0.33 0.33 0.33 RAT(L=128) 0.62 0.60 0.61 0.63 0.66 0.66 0.68 RAT(L=64) 0.67 0.64 0.66 0.70 0.73 0.74 1.03 RAT(L=16) 0.62 0.65 0.64 0.68 1.04 1.82 3.42 Attention 1.15 1.73 3.33 6.66 12.94 26.8 OOM $B=512$ RNN 0.35 0.34 0.33 0.33 0.34 0.34 0.33 RAT(L=128) 0.63 0.64 0.66 0.75 1.17 2.44 3.46 RAT(L=64) 0.70 0.68 0.74 1.16 2.41 3.44 6.69 RAT(L=16) 0.75 1.22 2.45 3.48 6.73 13.09 25.76 Attention 6.56 12.89 25.84 OOM OOM OOM OOM $B=1024$ RNN 0.36 0.36 0.35 0.35 0.35 0.36 0.35 RAT(L=128) 0.73 0.75 0.95 1.35 2.54 4.64 6.98 RAT(L=64) 0.74 0.94 1.35 2.48 4.67 7.02 13.45 RAT(L=16) 1.38 2.56 4.63 7.04 13.51 26.43 OOM Attention 13.04 25.70 OOM OOM OOM OOM OOM
@@
 As RAT also reduces cache memory usage, we report the maximum throughput of the full 1.3B model across different sequence lengths in Fig.
-5.
-For example, RAT(L=16) achieves an approximately 10.2×10.2$\times$10.2 × higher throughput than the baseline attention model at $T=4096$, and improves further to 15.6×15.6$\times$15.6 × at $T=16384$.
-Similarly, RAT(L=16)-SWA reaches 4.5×4.5$\times$4.5 ×, 8.0×8.0$\times$8.0 ×, and 13.9×13.9$\times$13.9 × higher throughput over attention at $T=4096$, $8192$, and $16384$, respectively.
-Figure 5: We measure the maximum throughput of the full 1.3B model for generating 1024 tokens under different prefilling lengths.
-For each total sequence length $T$, the prefilling length is set to $T-1024$.
-For example, $T=4096$ corresponds to a prefilling of 3072 tokens, while $T=8192$ and $T=16384$ correspond to 7168 and 15360 tokens, respectively.
-When the sequence length increases, the maximum throughput ratio between RAT and attention rises from 10.2×10.2$\times$10.2 × to 15.6×15.6$\times$15.6 ×, highlighting the strong efficiency advantage of RAT in long-context generation.
+4.
+For example, RAT(L=16) achieves an approximately $10.2\times$ higher throughput than the baseline attention model at $T=4096$, and improves further to $15.6\times$ at $T=16384$.
+Similarly, RAT(L=16)-SWA reaches $4.5\times$, $8.0\times$, and $13.9\times$ higher throughput over attention at $T=4096$, 81928192, and $16384$, respectively.
 Appendix C Accuracy
-C.1 Reversed design of RAT
+C.1 Supplementary ablation study
+All experiments here are conducted on the PG19 Book dataset with 200M-parameter models, and FLOPs are controlled for comparison.
+Reversed design of RAT
 Instead of employing recurrence within each chunk, we also explored a reversed design that applies attention inside the chunk, followed by recurrence to capture long-range dependencies.
 For this reversed design, standard RoPE is sufficient.
-However, with respect to parameter allocation, we found it crucial to retain the key vector in the attention module as a full tensor, as sharing it together with the query vector leads to a collapse into single-head attention.
-Thus, our final reversed design shares the query vector across attention heads, employs low-rank matrices for the projection used for output gate, and preserves full projections for both the forget gate and the key vector.
+Regarding the parameter allocation, we found it crucial to retain the key vector in the attention module as a full tensor, as sharing it together with the query vector leads to a collapse into single-head attention.
+Thus, the final reversed design shares the query vector across attention heads, employs low-rank matrices for the projection used for output gate, and preserves full projections for both the forget gate and the key vector.
 Results of the reversed design are put in Table 11.
 It can be observed that, under identical FLOP constraints, this design significantly underperforms RAT.
-For example, to match the FLOPs budget of $\mathcal{O}(T/64)$, it requires chunk sizes of 128, 256, and 512 for sequence lengths of 8192, 16384, and 32768, respectively.
-Most of the computational cost is spent within a small local window of 128 tokens, while excessive burden is placed on the inter-chunk recurrence, ultimately leading to high perplexity.
+We think this is because short-context dependencies are easier to capture and do not suffer from memory degradation, so using lightweight recurrence locally is more efficient.
+Long‑context dependencies are harder and may require direct retrieval of distant information, where attention is more suitable.
+For example, to match the FLOPs budget of $\mathcal{O}(T/64)$, the reversed design requires chunk sizes of 128, 256, and 512 for sequence lengths of 8192, 16384, and 32768, respectively.
+With this small local window size, the long-range dependencies are attributed to the recurrence, thus leading to high perplexity.
 Only when we reduce the number of chunks to 16 does the perplexity drop below 14.00.
 Therefore, while the reversed design also allows for interpolation between attention and RNN (reducing to attention when $L=T$, and to RNN when $L=1$), we choose to focus on RAT due to its more efficient utilization of FLOPs.
+Comparison to FLASH [gau]
+As pointed out in Sec.
+5, gau also uses the concept of chunking; however, we adopt a fundamentally different framework, both in the components involved and in how the two outputs are combined.
+Specifically, gau simply adds the outputs of inter-chunk and intra-chunk computations, whereas RAT organizes them hierarchically, applying recurrence over the key and value before inter-chunk attention.
+We argue that these differences lead to the weaker performance of gau, as shown in Table 13.
+First, as discussed above, RAT utilizes FLOPs more effectively than methods based on local attention, since its long-range dependencies must rely on memory-degrading components such as recurrence and linear attention.
+Second, we find that directly adding the outputs of softmax-based attention and linear attention introduces differences in output scale and potential representational conflicts, which may lower down the performance.
+Comparison to dilated attention
+We discuss the difference between RAT and the design that interleaves recurrence and dilated attention [longformer] with the sliding pattern.
+First, from an efficiency perspective, RAT has advantages over dilated attention in both training and inference.
+During training, dilated attention reduces parallelism by the dilation rate, since tokens within the same dilation span attend to different KV vectors.
+In contrast, RAT allows tokens to share the same chunk-level representation, except within their own chunk.
+During inference, the sliding dilation pattern causes a memory issue, since the KV cache of all tokens must be stored, whereas RAT reduces the cache size proportionally to the chunk size.
+Second, in terms of accuracy, to match the FLOPs, we set the dilation rate to 64 for RNN-Dilated attention, since dilated attention occurs in only half of the layers.
+It can be seen from Table 14 that dilated attention performs poorly.
+We observed significantly slower convergence at the beginning of training, likely due to its lack of global perception.
 Table 11: Perplexity results of the reversed design, where attention is applied within chunks and RNN is used across chunks.
@@
 Method PE Pretrain HellaSwag LAMBADA PIQA PPL acc_norm acc acc_norm Attention-SWA RoPE 7.61 57.1 48.0 72.7 RAT(L=16)-SWA RoPE 7.57 57.9 49.3 73.9 Attention-SWA NoPE 7.69 56.4 47.3 72.7 RAT(L=16)-SWA NoPE 7.63 56.8 47.8 73.7
+Table 13: Performance of different chunk-based designs with $T=16384$ and FLOPs $\mathcal{O}(T/64)$.
+Method Intra-chunk Inter-chunk Organization PPL RAT(L=64) Recurrence Attention Hierarchically 13.34 RAT-Reversed(C=64) Attention Recurrence Hierarchically 14.06 FLASH (C=128) Attention Linear attention Add together 14.4
+Table 14: Performance compared to dilated attention with sliding patterns under the same FLOPs.
+Method PPL RNN-Dilated attention 15.22 RAT(L=128) 13.44
 C.2 Length generalization
@@
 To study this, we consider SWA variants using either RoPE or NoPE for attention and RAT, following recent practices that interleave attention with local attention modules and apply positional encodings only in the local attention.
-This design has been adopted in practice and has shown promising results [51].
+This design has been adopted in practice and has shown promising results [cohere_llm].
 As shown in Fig.
-6, the RNN performs very steadily as the test sequence length increases.
+5, the RNN performs very steadily as the test sequence length increases.
 However, within the training context, it has the highest loss among all models.
@@
 While many techniques have been proposed to improve length extrapolation in attention models, they should also be considered for RAT, as RAT essentially redirects attention from every position to inter-chunk locations.
-For example, RoPE extrapolation in the attention module can be improved using methods from Peng et al.
-[52], bloc97 [53], and NoPE extrapolation has been studied in Wang et al.
-[54], which points out that NoPE still has a context length limit, though it performs better than RoPE.
+For example, RoPE extrapolation in the attention module can be improved using methods from yarn, rope_ntk, and NoPE extrapolation has been studied in nope_head_scale, which points out that NoPE still has a context length limit, though it performs better than RoPE.
 Their work links the failure to shifts in attention distributions and proposes tuning the temperature of attention heads to improve extrapolation.
 We believe such methods could also be applied to RAT, and leave further exploration of these techniques to future work.
-C.3 16K context window
-Table 11: Performance of 1.3B models with a 16K context window, including pretraining perplexity, commonsense reasoning accuracy (short-context), and SFT results on long-context tasks.
-For SFT, we focus on tasks that still have long context even with the 16K context window, such as QMSum and NarrativeQA3.
-For NarrativeQA3, we adopt the passage-only setting from the original paper instead of the summary+passage setup used in NarrativeQA2.
-Although passages are still truncated at 16K, most information are preserved, making passage-only a more meaningful and challenging evaluation setting.
-Model Throughput Pretrain Evaluation SFT
-token/sec Val. ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg. NarrativeQA3 QMSum
-RNN 130K 8.09 37.4 65.7 54.2 37.6 72.6 53.4 53.5 15.4 22.8
-RAT(L=16) 10376 7.52 36.5 64.8 55.8 45.6 71.5 53.5 54.6 21.6 24.2
-Attention 667 7.44 36.8 67.5 57.0 43.9 73.2 57.1 55.9 23.3 23.8
-Attention-SWA 1347 7.45 36.9 65.8 56.6 47.6 72.5 55.6 55.8 20.3 23.3
-RAT(L=16)-SWA 9244 7.42 38.5 67.5 57.6 47.7 73.3 57.0 56.9 22.6 24.6
-We also pretrain the models with a 16K context length.
-However, we observe that directly pretraining with such a long context window may lead to issues, as most web data is relatively short and needs to be concatenated by the separate token to fill the window.
-Due to the lack of naturally long-context data, the model may fail to learn true long-context capabilities.
-We believe this setting would benefit from more careful strategies, such as continued pretraining and balanced long-context data sources.
-Nonetheless, due to resource limitations, we directly pretrain the models on a 16K context length in this work, and hope that the results can reflect the ability of different architectures to handle longer contexts to some extent.
-Due to the short web data used during pretraining, we consider evaluating the models on short-context tasks, or applying SFT on long-context tasks that contain real long sequences.
-As shown in Table 11, overall, the performance of RAT(L=16) stays in between of RNN and attention, and RAT(L=16)-SWA performs the best.
-It achieves a 0.02 pretraining lower perplexity than attention, and improves by about 1 point on average in commonsense reasoning and 0.8 points on QMSum.
-However, it underperforms attention around 0.7 points in F1 score on the NarrativeQA3 task, unlike in the 4K context length setting.
-On the other hand, the maximum throughput improves significantly: from 4.5$\sim$ 4.5$\times$ 4.5 × compared to attention at the 4K context window to 13.9$\sim$ 13.9$\times$ 13.9 × at 16K.
-This suggests that, to pursue higher accuracy performance, we may consider increasing the local attention window size (e.g., from 1024 to 4096) under the 16K context length setting.
-C.4 Needle-in-Haystack (retrieval ability)
+C.3 Needle-in-Haystack (retrieval ability)
 To evaluate retrieval capabilities—an area where attention-based architectures are known to excel, and where RNNs typically fall short—we conduct experiments on the Needle-in-Haystack tasks.
-In these tasks, a “needle” (e.g., a magic number or UUID) is embedded within irrelevant passages, noisy text, or mixed with multiple key-value pairs.
+In these tasks, a "needle" (e.g., a magic number or UUID) is embedded within irrelevant passages, noisy text, or mixed with multiple key-value pairs.
 Each key-value pair is presented as a short identifier (the key) followed by its associated value, and the model is prompted to retrieve the correct value given a specific key.
-We adopt the RULER benchmark [55] and evaluate a range of Needle-in-Haystack task configurations.
+We adopt the RULER benchmark [ruler] and evaluate a range of Needle-in-Haystack task configurations.
 Specifically, task single_1 involves retrieving a specific number from a background filled with repeated noise.
@@
 And multikey_3 further incorporates the complex UUID format.
-In the multiquery and multivalue settings, the model is required to resolve multiple retrieval targets—either by answering several distinct queries, or by retrieving all values associated with a single key.
+In the multiquery and multivalue settings, the model is required to resolve multiple retrieval targets, either by answering several distinct queries or by retrieving all values associated with a single key.
 For detailed definitions of each task, we refer the reader to the benchmark documentation.
 During evaluation, we observed that models may fail to interpret certain prompts correctly.
-For instance, prompts like “A special magic number is hidden within the following text.
+For instance, prompts like "A special magic number is hidden within the following text.
 Make sure to memorize it.
-I will quiz you about the number afterwards.” can cause failures even in attention-based models, especially on the harder multikey tasks.
+I will quiz you about the number afterwards." can cause failures even in attention-based models, especially on the harder multikey tasks.
 To mitigate this, we apply a light, one-round supervised fine-tuning stage before evaluation to adapt the models to the instruction patterns.
-We generate 1,000 synthetic training samples for each of the 8 tasks, resulting in a total of 8,000 examples disjoint from the validation sets.
+We generate 1000 synthetic training samples for each of the 8 tasks, resulting in a total of 8000 examples disjoint from the validation sets.
 The models are trained on this dataset for one round and then are evaluated directly on all 8 tasks.
@@
 As a result, RAT naturally exhibits retrieval capabilities that lie between those of attention and RNNs.
-Table 12: Accuracy on the Needle-in-Haystack tasks with different configurations from the RULER benchmark [55], used to evaluate retrieval capabilities.
-Evaluation is conducted with a 4K context window, using exact match as the metric.
-Model single_1 single_2 single_3 multikey_1 multikey_2 multikey_3 multiquery multivalue
-RNN 99.4 99.8 0.0 2.2 0.0 0.0 5.8 5.7
-RAT(L=64) 100.0 99.8 55.6 96.4 62.2 1.0 86.3 89.1
-RAT(L=16) 99.6 100.0 94.6 99.6 99.6 82.6 91.2 94.8
-Attention 100.0 100.0 100.0 99.6 99.6 95.2 98.6 99.1
 Appendix D Broader Impacts